### v1

# arXiv:2603.01466v1[math.FA] 02 Mar 2026

Abstract.
Differently from the non-relativistic quantum mechanics, the violation of Bell inequalities in quantum field theory depends more on the structure of observable algebras (typically type III von Neumann algebras) rather than the choice of specific quantum states. Therefore, studying the violation of Bell inequalities based on the von Neumann algebraic framework often reveals information about the algebraic structure. In this paper, we employ three mutually-commuting von Neumann algebras to characterize quantum entanglement swapping networks, and establish Bell-like inequalities thereon, commonly referred to as bilocal inequalities. We investigate the algebraic structural conditions under which bilocal inequalities are satisfied or violated on the generated algebra of these three von Neumann algebras. Furthermore, the conditions for maximal violation of the inequalities can be utilized to infer the structural information of von Neumann algebras in reverse. Our results not only utilize the violation of bilocal inequalities to reveal the structural properties of von Neumann algebras, but can also be applied to quantum mechanics and quantum field theory.
Corresponding author
Key words and phrases: von Neumann algebra, Bell nonlocality, Bell inequality, Quantum network
1. Introduction
Bell nonlocality demonstrates that local measurements performed on one subsystem of a quantum state can instantaneously influence the measurement outcomes on another subsystem, regardless of the spatial separation between them [1, 2, 3]. Such nonlocal correlations can be detected through Bell inequalities, which serve as constraints that all local correlations must obey [4, 5, 6, 7, 8]. It has been demonstrated to offer quantum advantages in various device-independent quantum information tasks, including communication complexity [9], quantum key distribution [10, 11], randomness amplification [12, 13], and measurement-based quantum computation [14, 15].
Moreover, Bell nonlocality in quantum field theory (QFT) serves as a crucial bridge connecting quantum information science with fundamental physics [16, 17]. It provides a rigorous framework for reconciling quantum entanglement with relativistic causality, resolves conceptual challenges such as impossible measurements, and reveals how fundamental symmetries like parity violation affect quantum correlations [19, 18]. Furthermore, the tension between nonlocal quantum phenomena and the locality assumptions of general relativity offers profound insights for quantum gravity research, suggesting that spacetime itself may be an emergent property rather than a fundamental aspect of reality [20, 21, 22].
In QFT, the algebras assigned to touching spacetime regions are of type III von Neumann ones, so there are no pure normal states and no tensor-product structures [23, 24, 25]. This differs from the traditional quantum-mechanical setup, which is usually linked to type I von Neumann algebras and relies on the algebraic tensor product as its mathematical framework [27, 26, 28]. These are two distinct models, referred to respectively as the tensor product algebra model and the mutually-commuting von Neumann algebra model. In the tensor-product algebra model, a bipartite composite quantum system is described by the tensor product of two Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, denoted as $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}.$ Let $\mathcal{B}(\mathcal{H})$ denote the algebra of all bounded linear operators on $\mathcal{H}$, and $\mathcal{T}(\mathcal{H})$ the space of trace-class operators. Then there is a duality relation $\mathcal{B}(\mathcal{H})=\mathcal{T}(\mathcal{H})^{*}$. A quantum state $\rho$ is a positive operator with trace 1, and the set of states is denoted by $\mathcal{S}(\mathcal{H})=\{\rho\in\mathcal{T}(\mathcal{H})\mid\rho\geq 0,\ \mathrm{Tr}(\rho)=1\}$. The algebra of observables is the Jordan algebra of self-adjoint operators, denoted by $\mathcal{B}_{s}(\mathcal{H})=\{A\in\mathcal{B}(\mathcal{H})\mid A=A^{*}\}.$ However, it should be pointed out that relying solely on the above tensor-product model to discuss quantum information problems has two drawbacks [29, 30, 31]: 1) It fails to provide a universal framework for accurately describing phenomena in infinite-dimensional systems. 2) It is not suitable for studying traditional quantum physics such as quantum field theory, which requires the language of type III von Neumann algebras. In the mutually-commuting von Neumann algebra model, the algebra of observables of a quantum system is described by a von Neumann algebra $\mathcal{M}$, where $\mathcal{M}\subseteq\mathcal{B}(\mathcal{H})$, and $\mathcal{M}_{A},\mathcal{M}_{B}$ are two mutually commuting von Neumann subalgebras of $\mathcal{M}$. The notation $\mathcal{M}^{\prime\prime}$ denotes the double commutant of $\mathcal{M}$. However, in more general cases, a natural question arises: are the two models isomorphic? In other words, does $\mathcal{M}$ admit an isomorphism to $\mathcal{M}_{A}\otimes\mathcal{M}_{B}$? This question is intimately related to the famous Tsirelson problem, which is equivalent to the Connes embedding conjecture in von Neumann algebra theory [33, 32, 34]. The Tsirelson problem asks whether the sets of joint probability distributions of measurement outcomes are the same in the two models. In 2020, Ji and collaborators [34] designed a quantum interactive protocol based on the classical halting problem in computability theory, and proved that the answer to the Tsirelson problem is negative. This means that the tensor-product model algebra model and the mutually-commuting von Neumann algebra model are not equivalent. Therefore, the mutually-commuting von Neumann algebra model provides a more general framework, and many novel quantum phenomena fundamentally different from those in the tensor-product model have been discovered in this setting [34, 35, 36, 37].
Overall, Bell nonlocality is not merely a quantum peculiarity but a structural feature encoded in the classification of operator algebras, providing rigorous tools to quantify non-classical correlations in relativistic quantum systems [17]. Summers et al. introduces the maximal Bell correlation and proves that its value is bounded by $\sqrt{2}$ in mutually-commuting von Neumann algebra model, with equality attainable iff each subalgebra contains a copy of $\mathcal{M}_{2}({\mathbb{C}})$ [38]. Translating these bounds into the vacuum representation of algebraic quantum field theory, he shows that tangent wedge algebras are always maximally correlated, whereas strictly spacelike-separated wedges and decay exponentially with mass-governed distance [39, 40, 41, 42, 43]. These works expose a new algebraic invariant¡ªthe Bell correlation invariant¡ªthat distinguishes infinitely many isomorphism classes of mutually-commuting von Neumann algebra pairs and links maximal violation to the presence of the hyperfinite type-${\rm II}_{1}$ factor [17].
In contrast to entanglement originating from an individual source, quantum networks comprise numerous small-scale entangled states. Owing to the independence among distinct sources, the correlations emerging from quantum networks exhibit non-convex characteristics that transcend the polytopes associated with single-source entanglement [44, 45, 46, 47, 48, 49]. To date, Bell-type inequalities have been devised to certify nonlocal correlations across diverse network architectures, such as entanglement-swapping networks [44, 50], chain configurations [51, 52], star topologies [53, 54], polygon structures [55, 56, 57], tree-shaped networks [58, 59, 60], arbitrary acyclic networks [46, 47, 61], and arbitrary $k$-independent networks [62]. Alternative research directions examine the stronger forms of network nonlocality that surpass hybrid implementations involving classical variables and post-quantum resources [63, 64]. Furthermore, the interplay between graph theory and quantum correlations has attracted considerable attention [65, 66]. Nevertheless, limited progress has been made concerning the discrimination of correlations produced by different networks and the subsequent identification of underlying quantum topologies [67]. Interestingly, the notion of bi-locality in the mutually-commuting von Neumann algebra model has already been introduced by Ligthart et al [68], and Xu has addressed the inclusion problem between the tensor product algebra model and the mutually-commuting von Neumann algebra model in this setting [69].
In this paper we aim to establish bi-local inequalities within the mutually-commuting von Neumann algebra model and investigate how the degree of their violation is related to the structural properties of the algebras.
2. Ternary mutually-commuting von Neumann algebra models and entanglement swapping networks
Figure 1. An entanglement swapping network scenario with two sources. The connection between two parties represents the sharing of the physical system between them.
Quantum bilocal scenario. In non-relativistic quantum mechanics, the quantum entanglement swapping network (see Fig. 1) is a scenario of three parties consisting of Alice, Bob and Charles, and two sources $\rho_{AB}$, $\rho_{BC}$ shared between them. All the three parties can perform measurements (or inputs) on their system, labeled $x$, $y$ and $z$ for Alice, Bob and Charles. After measuring the three parties, they obtain outcomes denoted $a$, $b$ and $c$, respectively. Here, $x,y,z,a,b,c\in\{0,1\}$. Bob’s measurement $y$ might correspond to a joint measurement on the two systems that he receives from each source. The correlations between the measurement outcomes of the three parties are described by the joint probability distribution $p(abc|xyz)$. In this scenario, the joint probability distribution $p(abc|xyz)$ is bilocal if it can be written as
$p(abc|xyz)=\int\int d\lambda d\mu q_{1}(\lambda)q_{2}(\mu)p(a|x,\lambda)p(b|y,\lambda,\mu)p(c|z,\mu),$
where $\lambda$ ($\mu$) characterizes the joint hidden variable of the system produced by the source $\rho_{AB}$ ($\rho_{BC}$) [70, 71].
Let $A_{x},B_{y},C_{z}$ denote the measurements of Alice, Bob, and Charles with inputs $x$, $y$, $z$, respectively. Note that the middle party Bob always performs joint measurement. Then, the bilocal inequality can be written as:
(2.1) $\mathcal{S}\equiv\sqrt{|I|}+\sqrt{|J|}\leq 2,$
where
$I\equiv\sum_{x,z}\langle A_{x}B_{0}C_{z}\rangle=\langle(A_{0}+A_{1})B_{0}(C_{0}+C_{1})\rangle,$
$J\equiv\sum_{x,z}(-1)^{x+z}\langle A_{x}B_{1}C_{z}\rangle=\langle(A_{0}-A_{1})B_{1}(C_{0}-C_{1})\rangle$
and $\langle A_{x}B_{y}C_{z}\rangle=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}p(abc|xyz)$ [72].
In addition, one can choose different observables to construct different bilocal inequalities. For instance, the violation of the inequality
(2.2) $\mathcal{S}^{\prime}=|K|+|L|\leq 2$
can also be used to detect nonbilocal correlations, where $K=\langle A_{0}(B_{0}-B_{1})C_{0}\rangle$ and $L=\langle A_{1}(B_{0}+B_{1})C_{1}\rangle$ [73]. Of course, different inequalities exhibit varying capabilities in detecting nonlocality.
Mutually-commuting von Neumann algbera models. In QFT, the observables for Alice, Bob and Charles are associated with three mutually-commuting von Neumann algebras $\mathcal{M}_{A},\ \mathcal{M}_{B},\ \mathcal{M}_{C}$. In general, they are all required to be type III von Neumann algebra factors. Recall two von Neumann algebras satisfy Haag duality means that each algebra equals the commutant of the other. Haag duality encode quantum nonlocality intrinsically within their infinite-dimensional non-abelian structure through their type III factor property, which enforces that local algebras of spacelike separated regions are entangled with trivial relative commutants [74, 75]. Here, we do not restrict von Neumann algebras to III type factors, and generalize the Haag duality condition to mutual commutativity, proposing a von Neumann algebra model with mutual commutativity. Therefore, our model covers both scenarios of non-relativistic quantum mechanics and quantum field theory. The idea of this model is similar to that in Refs. [69, 68].
Definition 1. (Ternary mutually-commuting von Neumann algebra models) Let $\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C}$ be von Neumann subalgebras of $\mathcal{B}(H)$ over some Hilbert space $H$, and they are mutually commuting, i.e., $\mathcal{M}_{i}\subset\mathcal{M}_{j}^{\prime}$ with $i\neq j\in\{A,B,C\}$, where $\mathcal{M}_{j}^{\prime}$ is the communicator of $\mathcal{M}_{j}$. The generated von Neumann algebra
$\mathcal{M_{ABC}}=(\mathcal{M_{A}}\vee\mathcal{M_{B}}\vee\mathcal{M_{C}})^{\prime\prime}.$
We refer to the above model as the MCvNA model. When $\mathcal{M}_{ABC}=\mathcal{M}_{A}\otimes\mathcal{M}_{B}\otimes\mathcal{M}_{C}$, it is called the tensor product algebra model. Here, in order to correspond to the Hermitian property of quantum observables, these three algebras are always restricted to consist of self-adjoint elements.
We point out that the von Neumann algebra model defined above is intended for the study of algebraic structural properties; therefore, we have adopted the most general definition, which encompasses but does not correspond to models of quantum mechanics or quantum field theory.
We intend to use the above model of three mutually commuting von Neumann algebras to describe the entanglement swapping network in Fig. 1. We note that there is no correlation between parties $A$ and $C$ in the network. Mathematically, this independence can be described by the following formula. Let $\tau$ be a state in the dual space $\mathcal{M}_{ABC}^{*}$. In correspondence with the bilocal scenario, we always require that
() $\tau(ac)=\tau(a)\tau(c)$
for any $a\in\mathcal{M}_{A},c\in\mathcal{M}_{C}$. We call it as the independent condition (*). This assumption will be used throughout this paper.
3. Bilocal inequalities and their bounds
Following the assumptions in Section 2, the network correlation $\hat{p}=p(\alpha\beta\gamma|xyz)$ is defined as
$p(\alpha\beta\gamma|xyz)=\tau({A}_{\alpha|x}{B}_{\beta|y}{C}_{\gamma|z}),$
where the measurement operators ${A}_{\alpha|x}\in\mathcal{M}_{A},\ {B}_{\beta|y}\in\mathcal{M}_{B},\ {C}_{\gamma|z}\in\mathcal{M}_{C}$, $\{x,y,z\}$ and $\{\alpha,\beta,\gamma\}$ denote the inputs and outputs of measurement operators with $x,y,z,\alpha,\beta,\gamma\in\{0,1\}$. Let $a_{x}\in\mathcal{M}_{A}$, $b_{y}\in\mathcal{M}_{B}$, $c_{z}\in\mathcal{M}_{C}$ be operators such that $I\leq a_{x},b_{y},c_{z}\leq I$ for any $x,y,z\in\{0,1\}$, and $a_{x}=\sum_{\alpha=0}^{1}(-1)^{\alpha}A_{\alpha|x},b_{y}=\sum_{\beta=0}^{1}(-1)^{\beta}B_{\beta|y},c_{z}=\sum_{\gamma=0}^{1}(-1)^{\gamma}C_{\gamma|z}$. Then denote $I_{\tau}=\tau((a_{0}+a_{1})b_{0}(c_{0}+c_{1})),\ J_{\tau}=\tau((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))$. In the MCvNA model, analogous to Ineq. (2.1), we define
(3.1) $\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}.$
Analogous to Ineq. (2.2), under the same assumption, we have the following
(3.2) $\mathcal{S}_{\tau}^{\prime}=|K_{\tau}|+|L_{\tau}|.$
where, $K_{\tau}=\tau(a_{0}(b_{0}+b_{1})c_{0}),\ L_{\tau}=\tau(a_{1}(b_{0}-b_{1})c_{1})$.
Next, we rely on the abelianness of the algebra to determine the bounds for the bilocal inequality, respectively. We mention that all theorems in this section are based on the independence assumption (*).
The following conclusion indicates that in the mutually-commuting von Neumann algebra model, the maximum upper bound that the quantity (3.1) can achieve is $2\sqrt{2}$. This is analogous to the case in non-relativistic quantum mechanics, where the maximal violation of the bilocal value based on quantum resources is $2\sqrt{2}$.
Theorem 1. In the MCvNA model, $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2\sqrt{2}.$
Proof. According to Gelfand-Namark-Segal (GNS) construction, there is a -representation $\pi_{\tau}:\mathcal{M}_{ABC}\rightarrow H_{\tau}$ and a cyclic vector $\Omega\in H_{\tau}$ such that the set $\{\pi_{\tau}(d)\Omega:\ d\in\mathcal{M}_{ABC}\}$ is dense in $H_{\tau}$. It follows by applying the Cauchy-Schwarz inequality that
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}$
$=\sqrt{|\tau((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\tau((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}$
$\leq\sqrt{2}\sqrt{|\tau(b_{0}(a_{0}+a_{1})(c_{0}+c_{1}))|+|\tau(b_{1}(a_{0}-a_{1})(c_{0}-c_{1}))|}$
$=\sqrt{2}\sqrt{|\langle\pi_{\tau}(b_{0})\Omega,\pi_{\tau}((a_{0}+a_{1})(c_{0}+c_{1}))\Omega\rangle|+|\langle\pi_{\tau}(b_{1})\Omega,\pi_{\tau}((a_{0}-a_{1})(c_{0}-c_{1}))\Omega\rangle|}$
$\leq\sqrt{2}\sqrt{\sqrt{\|\pi_{\tau}(b_{0})\Omega\|\|\pi_{\tau}((a_{0}+a_{1})(c_{0}+c_{1}))\Omega\|}+\sqrt{\|\pi_{\tau}(b_{1})\Omega\|\|\pi_{\tau}((a_{0}-a_{1})(c_{0}-c_{1}))\Omega\|}}$
$\leq\sqrt{2}\sqrt{\sqrt{\tau((a_{0}+a_{1})^{2}(c_{0}+c_{1})^{2})}+\sqrt{\tau((a_{0}-a_{1})^{2}(c_{0}-c_{1})^{2})}}$
$\leq\sqrt{2}\sqrt{\sqrt{\tau(a_{0}+a_{1})^{2}+\tau(a_{0}-a_{1})^{2}}\sqrt{\tau(c_{0}+c_{1})^{2}+\tau(c_{0}-c_{1})^{2}}}$
$=\sqrt{2}\sqrt{\sqrt{2\tau(a_{0}^{2}+a_{1}^{2})}\sqrt{2\tau(c_{0}^{2}+c_{1}^{2})}}$
$\leq\sqrt{2}\sqrt{2\sqrt{2}\sqrt{2}}$
$=2\sqrt{2}.$
[…]
Below, we demonstrate how the value (3.1) depends on the abelianness of the algebras $\mathcal{M}_{A}$, $\mathcal{M}_{B}$ and $\mathcal{M}_{C}$. The results indicate that in entanglement-swapping networks, the abelianness of the three algebras plays distinct roles in setting the upper bound of the inequality to 2. This is not a simple generalization of the bipartite Bell scenario [17], where, with only two systems, Summers et al. showed that if one of these two algebras is abelian, the upper bound of the Bell inequality is 2.
Theorem 2. (1) If $\mathcal{M}_{A}$ and $\mathcal{M}_{C}$ are Abelian, then
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2.$
(2) If $\mathcal{M}_{B}$ is Abelian, then
$|K_{\tau}|+|L_{\tau}|\leq 2.$
Proof. To see (1), since $\mathcal{M}_{A}$ and $\mathcal{M}_{C}$ are Abelian, respectively, the eight elements
$a_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{0}a_{0})(1+\epsilon_{1}a_{1}),\ \ \ c_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{1}c_{0})(1+\epsilon_{1}c_{1})$
with $\epsilon_{0},\epsilon_{1}\in\{+,-\}$ are positive. By direct computation, one obtains that
$a_{0}+a_{1}=2(a_{++}-a_{--}),\ c_{0}+c_{1}=2(c_{++}-c_{--}),$
$a_{0}-a_{1}=2(a_{+-}-a_{-+}),\ c_{0}-c_{1}=2(c_{+-}-c_{-+}).$
So one obtains that
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\tau((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}$
$=2\left(\sqrt{|\tau((a_{++}-a_{--})b_{0}(c_{++}-c_{--}))|}+\sqrt{|\tau((a_{+-}-a_{-+})b_{1}(c_{+-}-c_{-+}))|}\right)$
$=2\left(\sqrt{|\tau(a_{++}b_{0}c_{++})-\tau(a_{++}b_{0}c_{--})-\tau(a_{--}b_{0}c_{++})+\tau(a_{--}b_{0}c_{--})|}\right$.
$\qquad\left.+\sqrt{|\tau(a_{+-}b_{1}c_{+-})-\tau(a_{+-}b_{1}c_{-+})-\tau(a_{-+}b_{1}c_{+-})+\tau(a_{-+}b_{1}c_{-+})|}\right)$
$\leq 2\left(\sqrt{|\tau(a_{++}c_{++})+\tau(a_{++}c_{--})+\tau(a_{--}c_{++})+\tau(a_{--}c_{--})|}\right$.
$\qquad\left.+\sqrt{|\tau(a_{+-}c_{+-})+\tau(a_{+-}c_{-+})+\tau(a_{-+}c_{+-})+\tau(a_{-+}c_{-+})|}\right)$
$=2\left(\sqrt{|\tau(a_{++}+a_{--})||\tau(c_{++}+c_{--})|}+\sqrt{|\tau(a_{+-}+a_{-+})||\tau(c_{+-}+c_{-+})|}\right)$
$\leq 2\sqrt{\tau(a_{++}+a_{--})+\tau(a_{+-}+a_{-+})}\sqrt{\tau(c_{++}+c_{--})+\tau(c_{+-}c_{-+})}$
$=2\sqrt{1}\sqrt{1}=2,$
where the first inequality follows the fact that $I\leq b_{i}\leq I$ and the order-preserving of state $\tau$, and the second inequality holds because of the Cauchy Schwarz inequality and the non-negativeness of $a_{\epsilon_{0}\epsilon_{1}},\ b_{\epsilon_{0}\epsilon_{1}}$.
To prove (2), since $\mathcal{M}_{B}$ is Abelian, the four elements
$b_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{0}a_{0})(1+\epsilon_{1}a_{1}),$
with $\epsilon_{0},\epsilon_{1}\in\{+,-\}$ are positive. It follows from a direct computation that
$b_{0}+b_{1}=2(b_{++}-b_{--}),\ b_{0}-b_{1}=2(b_{+-}-b_{-+}).$
So one obtains that
$\tau[a_{0}(b_{0}+b_{1})c_{0}]+\tau[a_{1}(b_{0}-b_{1})c_{1}]$
$=$ $2(\tau[a_{0}(b_{++}-b_{--})c_{0}]+\tau[a_{1}(b_{+-}-b_{-+})c_{1}])$
$\leq$ $2(\tau[b_{++}]+\tau[b_{--}]+\tau[b_{+-}]+\tau[b_{-+}])=2$
[…]
Remark. The above theorem illustrates a phenomenon: the violations of the two Bell-type inequalities $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2$ and $|K_{\tau}|+|L_{\tau}|\leq 2$ can indicate the abelianness of the underlying algebras. Conversely, applying the theorem, we can deduce from the violation of an inequality with an upper bound of 2 that the three von Neumann algebras are non-abelian. Here, one also needs to consider the chosen observables and the specific state.
If we wish to set aside the influence of the algebraic structure and consider separately the impact of the choice of observables and the state on values of (3.1) and (3.2), then the following conclusion can be regarded as a generalization of the results in the non-classical mechanical framework. For the sake of clarity, we introduce a tensor product structure into the three-algebra MCvNA model. Let $\mathcal{M}_{B}=\mathcal{M}_{B}^{L}\otimes\mathcal{M}_{B}^{R}$. We call the state $\tau\in\mathcal{M}_{ABC}^{*}$ is separable on $(\mathcal{M}_{A}\vee\mathcal{M}_{B}^{L})^{\prime\prime}$ if there are $\xi_{\alpha}\in\mathcal{M}_{A}^{*},\eta_{\alpha}\in(\mathcal{M}_{B}^{L})^{*}$ such that $\tau(ab^{L})=\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(a)\eta_{\alpha}(b^{L})$ with $\sum_{\alpha}\lambda_{\alpha}=1$.
Theorem 3. $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2$ for any state $\tau$ that is separable on $(\mathcal{M}_{A}\vee\mathcal{M}_{B}^{L})^{\prime\prime}$ and $(\mathcal{M}_{B}^{R}\vee\mathcal{M}_{C})^{\prime\prime}$.
Proof. It follows from separability assumption of $\tau$ that
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\tau((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}$
$=\sqrt{|\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(a_{0}+a_{1})\eta_{\alpha}(b_{0}^{L})||\sum_{\beta}\mu_{\beta}M_{\beta}(b_{0}^{R})N_{\beta}(c_{0}+c_{1})|}$
$\hskip 18.49988pt+\sqrt{|\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(a_{0}-a_{1})\eta_{\alpha}(b_{1}^{R})||\sum_{\beta}\mu_{\beta}M_{\beta}(b_{1}^{R})N_{\beta}(c_{0}-c_{1})|}$
$\leq\sqrt{(\sum_{\alpha}\lambda_{\alpha}|\xi_{\alpha}(a_{0}+a_{1})|)(\sum_{\beta}\mu_{\beta}|N_{\beta}(c_{0}+c_{1})|)}+\sqrt{(\sum_{\alpha}\lambda_{\alpha}|\xi_{\alpha}(a_{0}-a_{1})|)(\sum_{\beta}\mu_{\beta}|N_{\beta}(c_{0}-c_{1})|)}$
$\leq\sqrt{\sum_{\alpha}\lambda_{\alpha}(|\xi_{\alpha}(a_{0}+a_{1})|+|\xi_{\alpha}(a_{0}-a_{1})|)}\sqrt{\sum_{\beta}\mu_{\beta}(|N_{\beta}(c_{0}+c_{1})|+|N_{\beta}(c_{0}-c_{1})|)}$
$=2\sqrt{\sum_{\alpha}\lambda_{\alpha}\max\{\xi_{\alpha}(a_{0}),\xi_{\alpha}(a_{1})\}}\sqrt{\sum_{\beta}\mu_{\beta}\max\{N_{\beta}(c_{0}),N_{\beta}(c_{1})\}}$
$\leq 2.$
[…]
From the above conclusion, we see that whether a bilocal inequality can be violated depends on the choice of observables and the state, and even more so on the structure of the algebra. In the following, we define a quantity
$\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})=\sup_{\{a_{x},b_{y},c_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}).$
Combining Theorems 1, 2 and 3, one naturally obtains the following corollary. A similar conclusion can be drawn for the quantity (3.2).
Corollary 4.
(1) $2\leq\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})\leq 2\sqrt{2}$.
(2) If $\mathcal{M}_{A}$ and $\mathcal{M}_{C}$ are Abelian, then $\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})=2$.
(3) $\phi\in(\mathcal{M}_{A}\vee\mathcal{M}_{B_{1}}\vee\mathcal{M}_{B_{2}}\vee\mathcal{M}_{C})^{*}$ is convex combination of separable states, then $\mathcal{S}(\phi,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})=2$.
(4) $|\mathcal{S}(\phi,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})-\mathcal{S}(\psi,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})|\leq k\sqrt{\|\phi-\psi\|}$, so, in particular, $\sup_{\{a_{x},b_{y},c_{z}\}}(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|})$ is norm continuous.
Proof. To show (1). Since $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $a_{0}=a_{1}=b_{0}=c_{0}=c_{1}=I$. And combinig the proof of Theorem 2, we obtain (1).
(2) holds by facts $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $a_{0}=a_{1}=b_{0}=c_{0}=c_{1}=I$ and Theorem 2.
(3) holds by facts $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $a_{0}=a_{1}=b_{0}=c_{0}=c_{1}=I$ and Theorem 3.
To prove (4). According to the representations of $I_{\tau}$ and $J_{\tau}$, and the facts that $|\sup x-\sup y|\leq\sup|x-y|$ and triangle inequality, one can obtain that
$\left|\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|}\right)-\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{|I_{\psi}|}+\sqrt{|J_{\psi}|}\right)\right|$
$=\left|\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{|\phi((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\phi((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}\right)\right$.
$\qquad\left.-\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{|\psi((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\psi((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}\right)\right|$
$\leq\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left|\sqrt{|\phi((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\phi((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}\right$.
$\qquad\left.-\sqrt{|\psi((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}-\sqrt{|\psi((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}\right|$
$\leq\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\left|\sqrt{|\phi((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}-\sqrt{|\psi((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}\right|\right$.
$\qquad\left.+\left|\sqrt{|\phi((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}-\sqrt{|\psi((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}\right|\right)$
$\leq\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{|(\phi-\psi)((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|(\phi-\psi)((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}\right)$
$\leq\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{\|\phi-\psi\|\,\|(a_{0}+a_{1})b_{0}(c_{0}+c_{1})\|}+\sqrt{\|\phi-\psi\|\,\|(a_{0}-a_{1})b_{1}(c_{0}-c_{1})\|}\right)$
$\leq k\sqrt{\|\phi-\psi\|},$
[…]
4. Maximal violation of bilocal inequalities and algebraic structures
In this section, we aim to point out that the violation of bilocal inequalities, in particular the maximal violation, can reflect the structural properties of the algebra. Here, violation refers to exceeding the upper bound of 2 for the inequality, and maximal violation refers to attaining $2\sqrt{2}$.
In the following theorem, we attempt to analyze the conditions for maximal violation.
Theorem 5. If $\tau$ is faithful, then
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$
if and only if $\tau(a_{i}^{2}a)=\tau(a),\ \tau(b_{i}^{2}b)=\tau(b),\ \tau(c_{i}^{2}c)=\tau(c)$, and $\tau[(a_{0}a_{1}+a_{1}a_{0})a]=0,\ \tau[(c_{0}c_{1}+c_{1}c_{0})c]=0$ for any $a\in\mathcal{M}_{A},b\in\mathcal{M}_{B},c\in\mathcal{M}_{C}$.
Remark. This theorem states that when the inequality is maximally violated, $I,\ a_{0},\ a_{1}$ and $\frac{i}{2}[a_{0},a_{1}]$ form a realization of the Pauli spin matrices in $\mathcal{M}_{A}$ and generate a copy of $M_{2}(\mathbb{C})$ in $\mathcal{M}_{A}$. And the same holds for $I,\ c_{0},\ c_{1}$ and $\frac{i}{2}[c_{0},c_{1}]$ in $\mathcal{M}_{C}$. Moreover, $\tau(a_{0}a_{1}+a_{1}a_{0})=0,\ \tau(c_{0}c_{1}+c_{1}c_{0})=0.$
Proof. If $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$, it follows from the proof of Theorem 1 that for any $t\in[0,1]$, these equalities hold:
(4.1) $|\tau(b_{0}(a_{0}+a_{1})(c_{0}+c_{1}))|=|\tau(b_{1}(a_{0}-a_{1})(c_{0}-c_{1}))|,$
(4.2) $\pi_{\tau}(b_{0})\Omega=k_{0}\pi_{\tau}[(a_{0}+a_{1})(c_{0}+c_{1})]\Omega,$
(4.3) $\pi_{\tau}(b_{1})\Omega=k_{1}\pi_{\tau}[(a_{0}-a_{1})(c_{0}-c_{1})]\Omega,$
(4.4) $\|\pi_{\tau}(b_{0})\Omega\|=\|\pi_{\tau}(b_{1})\Omega\|=1,$
(4.5) $\tau[(a_{0}+a_{1})^{2}]=t\tau[(c_{0}+c_{1})^{2}],$
(4.6) $\tau[(a_{0}-a_{1})^{2}]=t\tau[(c_{0}-c_{1})^{2}],$
(4.7) $a_{0}^{2}+a_{1}^{2}=2I,c_{0}^{2}+c_{1}^{2}=2I.$
From the (4.7), one gets $a_{i}^{2}=I,\ c_{i}^{2}=I\ (i=0,1)$ because $I\leq a_{i},c_{i}\leq I$, therefore $\tau(a_{i}^{2}a)=\tau(a)\ \tau(c_{i}^{2}c)=\tau(c)$ for any $a\in\mathcal{M}_{A},\ b\in\mathcal{M}_{B},\ c\in\mathcal{M}_{C}$.
Now let us show the proof for $\tau(b_{i}^{2}b)=\tau(b),\ \tau[(a_{0}a_{1}+a_{1}a_{0})a]=0$, and $\tau[(c_{0}c_{1}+c_{1}c_{0})c]=0$ for any $a\in\mathcal{M}_{A},\ b\in\mathcal{M}_{B},\ c\in\mathcal{M}_{C}$, which implies $b_{i}^{2}=I,\ a_{0}a_{1}+a_{1}a_{0}=0,$ and $c_{0}c_{1}+c_{1}c_{0}=0$. According to the Eqs. (4.5)-(4.7), i.e.,
$\left\{\begin{array}[]{c}\tau(a_{0}^{2}+a_{1}^{2}+a_{0}a_{1}+a_{1}a_{0})=t\tau(c_{0}^{2}+c_{1}^{2}+c_{0}c_{1}+c_{1}c_{0})\\ \tau(a_{0}^{2}+a_{1}^{2}-a_{0}a_{1}-a_{1}a_{0})=t\tau(c_{0}^{2}+c_{1}^{2}-c_{0}c_{1}-c_{1}c_{0})\\ a_{0}^{2}+a_{1}^{2}=2I,\ c_{0}^{2}+c_{1}^{2}=2I,\end{array}\right$.
we can get
$t=1,\ \tau(a_{0}a_{1}+a_{1}a_{0})=\tau(c_{0}c_{1}+c_{1}c_{0}).$
Then combing the conditions (4.2), (4.3) and condition (4.4) one can get
$k_{0}^{2}\tau[(a_{0}+a_{1})^{2}(c_{0}+c_{1})^{2}]=1,\ k_{1}^{2}\tau[(a_{0}-a_{1})^{2}(c_{0}-c_{1})^{2}]=1,$
i.e.,
(4.8) $|k_{0}|=\frac{1}{2+\tau(X)},\ \ \ |k_{1}|=\frac{1}{2-\tau(X)}$
because of $\tau(X)=\tau(Y)$, where $X=a_{0}a_{1}+a_{1}a_{0},\ Y=c_{0}c_{1}+c_{1}c_{0}$. Then according to Eqs. (4.2), (4.3), and (4.1),
$|\tau[b_{0}(a_{0}+a_{1})(c_{0}+c_{1})]|=|k_{0}||\tau[(a_{0}+a_{1})^{2}(c_{0}+c_{1})^{2}]|,$
$|\tau[b_{1}(a_{0}-a_{1})(c_{0}-c_{1})]|=|k_{1}|\tau[(a_{0}-a_{1})^{2}(c_{0}-c_{1})^{2}]|.$
Substituting (4.8) and (4.1) to the above equations, we have
$\frac{1}{2+\tau(X)}(2+\tau(X))^{2}=\frac{1}{2-\tau(X)}(2-\tau(X))^{2}$
deriving that $\tau(X)=\tau(Y)=0$. It follows from $\tau(X)=0=\tau(Y)$ and $\|\pi_{\tau}(b_{i})\Omega\|=1$ that
$k_{0}^{2}=k_{1}^{2}=\frac{1}{4}$
from the Eq. (4.8). Since
$|\langle\Omega,\pi_{\tau}(b_{0}^{2})\Omega\rangle|$ $=$ $|\frac{1}{4}\langle\Omega,\pi_{\tau}[(a_{0}+a_{1})^{2}(c_{0}+c_{1})^{2}]\Omega\rangle|$
$\leq$ $\frac{1}{4}\sqrt{\|\pi_{\tau}[(a_{0}+a_{1})(c_{0}+c_{1})]\Omega\|}\sqrt{\|\pi_{\tau}[(a_{0}+a_{1})(c_{0}+c_{1})]\Omega\|}$
$=$ $\frac{1}{4}\tau[(a_{0}+a_{1})^{2}]\tau[(c_{0}+c_{1})^{2}]=1,$
and $\tau(b_{0}^{2})=|\langle\Omega,\pi_{\tau}(b_{0}^{2})\Omega\rangle|=1$, this implies that $\pi_{\tau}[(b_{0}^{2})]\Omega=\Omega$, so
$\tau(b_{0}^{2}b)=\langle\Omega,\pi_{\tau}(b_{0}^{2})\pi_{\tau}(b)\Omega\rangle=\langle\pi_{\tau}(b_{0}^{2})\Omega,\pi_{\tau}(b)\Omega\rangle=\langle\Omega,\pi_{\tau}(b)\Omega\rangle=\tau(b).$
Similarly $\tau(b_{1}^{2}b)=\tau(b)$ for any $b\in\mathcal{M}_{B}.$ Furthermore, it follows from Eq. (4.2) and $\pi_{\tau}[(b_{0}^{2})]\Omega=\Omega$ that
$\pi_{\tau}(2X+2Y+XY)\Omega=0.$
So $\tau[(2X+2Y+XY)^{2}]=0,$ implying that
$\tau(X^{2})=\tau(Y^{2})=0.$
Combining the faithfulness, non-negativity of the state $\tau$ and the self-adjointness of X,YX,\ Y. Then $X=Y=0$, i.e.,
$a_{0}a_{1}+a_{1}a_{0}=c_{0}c_{1}+c_{1}c_{0}=0,$
and implies that $\tau[(a_{0}a_{1}+a_{1}a_{0})a]=\tau[(c_{0}c_{1}+c_{1}c_{0})c]=0$ .
[…]
The above Theorem 5 indicates that if the bilocal inequality is maximally violated, then the underlying algebra locally contains an $M_{2}(\mathbb{C})$ factor. However, this does not lead to further global structural properties of the algebra. Below, we aim to further explore the relationship between the maximal violation of the inequality and the global properties of the algebra.
Theorem 6. Let $\mathcal{M}_{ABC}$ be a von Neumann algebra on a separable Hilbert space $H$ with cyclic and separating vector. If there are hyperfinite type II1 factors $\mathcal{R}_{A}$ and $\mathcal{R}_{C}$ such that $\mathcal{M}_{A}\simeq\mathcal{M}_{A}\otimes\mathcal{R}_{A}$ and $\mathcal{M}_{C}\simeq\mathcal{M}_{C}\otimes\mathcal{R}_{C}$, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$.
Proof. We first deal with the special case: $\mathcal{M}_{A}=\mathcal{R}_{A}$ and $\mathcal{M}_{C}=\mathcal{R}_{C}$. Let $\widetilde{\mathcal{H}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{H}_{\alpha},\Omega_{\alpha})$ be the incomplete tensor product of $\mathcal{H}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Omega_{\alpha}=\frac{1}{\sqrt{2}}(\Phi_{1}\otimes\Phi_{1}+\Phi_{2}\otimes\Phi_{2})\equiv\Omega\text{ for some basis }\{\Phi_{1},\Phi_{2}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then by the construction of Araki and Woods [76], $\mathcal{R}_{A}\simeq\mathcal{R}(\mathcal{H}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Omega_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{H}}$. Similarly, let $\widetilde{\mathcal{K}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{K}_{\alpha},\Lambda_{\alpha})$ be the incomplete tensor product of $\mathcal{K}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Lambda_{\alpha}=\frac{1}{\sqrt{2}}(\Psi_{3}\otimes\Psi_{3}+\Psi_{4}\otimes\Psi_{4})\equiv\Lambda\text{ for some basis }\{\Psi_{3},\Psi_{4}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then $\mathcal{R}_{C}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{K}}$. We can obtain that there is $\mathcal{R}_{C}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$. Now from Theorem 5, we know that $I,\ a_{0},\ a_{1}$ and $a_{2}=-\frac{i}{2}[a_{0},a_{1}]$ form a realization of the Pauli spin matrices on $\mathcal{H}_{\alpha}$, and the same holds for $I,\ c_{0},\ c_{1}$ and $c_{2}=-\frac{i}{2}[c_{0},c_{1}]$ in $\mathcal{K}_{\alpha}$, and $b_{0},b_{1}$ with $b_{i}^{2}=I$ such that the quantity (3.1) arrives at $2\sqrt{2}$. Applying the similar technologies of Eqs. (6)-(8) of Ref.[72], one can obtain that
(4.9) $\sqrt{|\langle\Omega\otimes\Lambda,[(a_{0}+a_{1})b_{0}(c_{0}+c_{1})]\Omega\otimes\Lambda\rangle|}=\sqrt{2},$
(4.10) $\sqrt{|\langle\Omega\otimes\Lambda,[(a_{0}-a_{1})b_{1}(c_{0}-c_{1})]\Omega\otimes\Lambda\rangle|}=\sqrt{2}$
[…]
$\widetilde{T}_{\alpha}=(\widetilde{a}_{0,\alpha}+\widetilde{a}_{1,\alpha})\widetilde{b}_{0,\alpha}(\widetilde{c}_{0,\alpha}+\widetilde{c}_{1,\alpha}),$
$\widetilde{S}_{\alpha}=(\widetilde{a}_{0,\alpha}-\widetilde{a}_{1,\alpha})\widetilde{b}_{1,\alpha}(\widetilde{c}_{0,\alpha}-\widetilde{c}_{1,\alpha}).$
and furthermore $\Phi=\bigotimes_{\alpha\in\mathbb{N}}\Phi_{\alpha},\Psi=\bigotimes_{\alpha\in\mathbb{N}}\Psi_{\alpha},$ with $\Phi_{\alpha}=\Omega_{\alpha}\otimes\Omega_{\alpha}$ and $\Psi_{\alpha}=\Lambda_{\alpha}\otimes\Lambda_{\alpha}$ for almost all $\alpha$. By definition of the incomplete infinite tensor product, such vectors are total in $\widetilde{\mathcal{H}}\otimes\widetilde{\mathcal{K}}$. But one has
$\sqrt{\langle\Phi,\widetilde{T}_{\alpha}\Psi\rangle}=\langle\Phi,\Psi\rangle\cdot\sqrt{2},\sqrt{\langle\Phi,\widetilde{S}_{\alpha}\Psi\rangle}=\langle\Phi,\Psi\rangle\cdot\sqrt{2}.$
for all $\alpha$ larger than some $\alpha_{0}$ (depending on $\Phi$ and $\Psi$).
[…]
It is mentioned that the results in Theorem 6 can be applied into the quantum field theory. In the algebraic framework of the QFT, wedge algebras (von Neumann algebras associated to specific unbounded regions in Minkowski spacetime, such as $\mathcal{O}$ associated to tangent regions are typically type III1 factors (a classification indicating the algebra has no trace and no pure normal states, reflecting the infinite degrees of freedom of QFT). By the classical result of A. Connes, all injective (i.e., amenable or hyperfinite, a property related to physical phase space compactness) infinite factors (excepting some type III0) are strongly stable (meaning they can absorb a hyperfinite II1 factor: $\mathcal{M}\simeq\mathcal{M}\otimes\mathcal{R}_{1}$). So the assumption of Theorem 6 is satisfied.
Corollary 7. Let $\mathcal{O}_{i}$ $(i=1,2,3)$ be wedge-shaped regions, $\mathcal{A}(\mathcal{O}_{i})$ the wedge algebra on a separable Hilbert space $H$ with cyclic and separating vector. If each $\mathcal{A}(\mathcal{O}_{i})$ is III1 factor, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$ on $(\mathcal{A}(\mathcal{O}_{1})\vee\mathcal{A}(\mathcal{O}_{2})\vee\mathcal{A}(\mathcal{O}_{3}))^{\prime\prime}$.
Acknowledgement. All authors contributed equally to this work.