### v2

# Violation of Quantum Bilocal Inequalities on Mutually-Commuting von Neumann Algebra Models

Abstract.
In contrast to the non-relativistic quantum mechanics, the violation of Bell inequalities in quantum field theory depends more on the structure of observable algebras (typically type III von Neumann algebras) rather than the choice of specific quantum states. Therefore, studying the violation of Bell inequalities based on the von Neumann algebraic framework can be applied to addressing questions in quantum field theory, while also providing insights into the structural properties of von Neumann algebras. In this paper, we employ three mutually-commuting von Neumann algebras to characterize quantum entanglement swapping networks, and establish Bell-like inequalities thereon, commonly referred to as bilocal inequalities. We investigate the algebraic structural conditions under which bilocal inequalities are satisfied or violated on the generated algebra of these three von Neumann algebras. Furthermore, the conditions for maximal violation of the inequalities can be utilized to infer the structural information of von Neumann algebras in reverse.
Corresponding author
Key words and phrases: von Neumann algebra, Bell nonlocality, Bell inequality, Quantum network
1. Introduction
In non-relativistic quantum mechanics, Bell nonlocality demonstrates that local measurements performed on one subsystem of a quantum state can instantaneously influence the measurement outcomes on another subsystem, regardless of the spatial separation between them [1, 2, 3]. Such nonlocal correlations can be detected through Bell inequalities, which serve as constraints that all local correlations must obey [4, 5, 6, 7, 8]. It has been demonstrated to offer quantum advantages in various device-independent quantum information tasks, including communication complexity [9], quantum key distribution [10, 11], randomness amplification [12, 13], and measurement-based quantum computation [14, 15]. However, it should be pointed out that relying solely on the non-relativistic quantum mechanics to discuss quantum information problems has two drawbacks [16, 17, 18]: 1) it fails to provide a universal framework for accurately describing phenomena in systems with infinite degrees of freedom; 2) it is not suitable for studying traditional quantum physics such as quantum field theory, which requires the language of type III von Neumann algebras. Research on quantum information problems on von Neumann algebras has received significant attention and yielded many meaningful results [19, 20, 21, 22, 23, 24, 25].
In the 1980s early, mathematician Summers et al. introduce firstly the maximal violation of Bell inequalities and prove that its value is bounded by $2\sqrt${2} in the von Neumann algebra model, with equality attainable iff each subalgebra contains a copy of $\mathcal{M}_{2}({\mathbb{C}})$ [26]. This shows that Bell nonlocality is not merely a quantum peculiarity but a structural feature encoded in the classification of operator algebras, providing rigorous tools to quantify non-classical correlations in relativistic quantum systems [27]. Translating these bounds into the vacuum representation of algebraic quantum field theory, they show that tangent wedge algebras are always maximally correlated, whereas strictly spacelike-separated wedges decay exponentially with mass-governed distance [28, 29, 30, 31, 32]. These works reveal a novel algebraic invariant, termed the Bell correlation invariant, which distinguishes infinitely many isomorphism classes of pairs of mutually commuting von Neumann algebras and links the maximal violation to the occurrence of the hyperfinite type $\rm II_{1}$ factor [27]. This is a pioneering work to make Bell nonlocality in quantum field theory (QFT) serve as a crucial bridge connecting quantum information science with fundamental physics [27, 33]. It provides a rigorous framework for reconciling quantum entanglement with relativistic causality, resolves conceptual challenges such as impossible measurements, and reveals how fundamental symmetries like parity violation affect quantum correlations [34, 35, 36, 37, 38].
In contrast to entanglement originating from an individual source, quantum networks comprise numerous small-scale entangled states. Owing to the independence among distinct sources, the correlations emerging from quantum networks exhibit non-convex characteristics that transcend the polytopes associated with single-source entanglement [39, 40, 41, 42, 43, 44]. To date, Bell-like inequalities in the non-relativistic quantum mechanics have been devised to certify nonlocal correlations across diverse network architectures, such as entanglement-swapping networks [39, 45], chain configurations [46, 47], star topologies [48, 49], polygon structures [50, 51, 52], tree-shaped networks [53, 54, 55], arbitrary acyclic networks [41, 42, 56], and arbitrary $k$-independent networks [57]. Alternative research directions examine the stronger forms of network nonlocality that surpass hybrid implementations involving classical variables and post-quantum resources [58, 59]. Nevertheless, limited progress has been made concerning the discrimination of correlations produced by different networks and the subsequent identification of underlying quantum network topologies [60]. Recently, the notion of bi-locality in an entanglement swapping network based on the mutually-commuting von Neumann algebra model has already been introduced by Ligthart et al. [61], and Xu has addressed the inclusion problem between the tensor product algebra model and the mutually-commuting von Neumann algebra model in this setting [62]. However, Bell-like inequalities for such von Neumann algebraic models have not yet been established. In this paper, we aim to establish bi-local inequalities within the mutually-commuting von Neumann algebra model and investigate how the degree of their violation is related to the structural properties of the algebras.
2. Mutually-commuting von Neumann algebra models, Tsirelson problem, and Connes’ embedding conjecture
In QFT, the algebras assigned to touching spacetime regions are of type III von Neumann ones, so there are no pure normal states and no tensor-product structures [63, 64, 65]. This differs from the non-relativistic quantum-mechanical setup, which is usually linked to type I von Neumann algebras and relies on the algebraic tensor product as its mathematical framework [66, 67, 68]. These are two distinct models, referred to respectively as the tensor product algebra model and the mutually-commuting von Neumann algebra model.
In the tensor-product algebra model, a bipartite composite quantum system is described by the tensor product of two Hilbert spaces $\mathcal{H}_{\mathcal{A}}$ and $\mathcal{H}_{\mathcal{B}}$, denoted as $\mathcal{H}=\mathcal{H}_{\mathcal{A}}\otimes\mathcal{H}_{\mathcal{B}}.$ Let $\mathcal{B}(\mathcal{H})$ denote the algebra of all bounded linear operators on $\mathcal{H}$, and $\mathcal{T}(\mathcal{H})$ the space of trace-class operators. Then there is a duality relation $\mathcal{B}(\mathcal{H})=\mathcal{T}(\mathcal{H})^{*}$. A quantum state $\rho$ is a positive semidefinite operator with trace 1, and the set of states is denoted by $\mathcal{S}(\mathcal{H})=\{\rho\in\mathcal{T}(\mathcal{H})\mid\rho\geq 0,\ \mathrm{Tr}(\rho)=1\}$. The algebra of observables is the Jordan algebra of self-adjoint operators, denoted by $\mathcal{B}_{s}(\mathcal{H})=\{A\in\mathcal{B}(\mathcal{H})\mid A=A^{*}\}.$ In the mutually-commuting von Neumann algebra model, the algebra of observables of a quantum system is described by a von Neumann algebra $\mathcal{M}$, with $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{B}}$ being two mutually commuting von Neumann subalgebras of $\mathcal{M}$ such that $(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}})^{\prime\prime}=\mathcal{M}$. Here, $\mathcal{M}^{\prime\prime}$ denotes the double commutant of $\mathcal{M}$.
However, in more general cases, a natural question arises: are the two models isomorphic? In other words, does $\mathcal{M}$ admit an isomorphism to $\mathcal{M}_{\mathcal{A}}\otimes\mathcal{M}_{\mathcal{B}}$? This question is intimately related to the famous Tsirelson problem, which is equivalent to the Connes embedding conjecture in von Neumann algebra theory [69, 70, 71]. The Tsirelson problem asks whether the sets of joint probability distributions of measurement outcomes are the same in the two models. In 2020, Ji and collaborators [71] designed a quantum interactive protocol based on the classical halting problem in computability theory, and proved that the answer to the Tsirelson problem is negative. This means that the tensor-product algebra model and the mutually-commuting von Neumann algebra model are not equivalent. Therefore, the mutually-commuting von Neumann algebra model provides a more general framework, and many novel quantum phenomena fundamentally different from those in the tensor-product model have been discovered in this setting [71, 72, 73, 74].
3. Ternary mutually-commuting von Neumann algebra models and entanglement swapping networks
Figure 1. An entanglement swapping network scenario with two sources. The connection between two parties represents the sharing of the physical system between them.
Quantum bilocal scenario. In non-relativistic quantum mechanics, the quantum entanglement swapping network (see Fig. 1) is a scenario of three parties consisting of Alice, Bob and Charles, and two sources $\rho_{AB}$, $\rho_{BC}$ shared between them. The inputs and outputs of the measurements performed by the three parties are denoted as $x,y,z$ and $a,b,c$, respectively, where $x,y,z,a,b,c\in\{0,1\}$. Assume that each party performs binary-input and binary-output measurements, with the observables for Alice, Bob, and Charles denoted as $A_{x}$, $B_{y}$ and $C_{z}$, respectively. Here it is required that the spectra of operators $A_{x},B_{y},C_{Z}$ are all {-1,1}, implying that $I\leq A_{x},B_{y},C_{z}\leq I$. The correlations between the measurement outcomes of the three parties are described by the joint probability distribution $p(abc|xyz)$. In this scenario, $p(abc|xyz)$ is said to be bilocal if it can be written as
$p(abc|xyz)=\int\int d\lambda d\mu p_{1}(\lambda)p_{2}(\mu)p(a|x,\lambda)p(b|y,\lambda,\mu)p(c|z,\mu),$
where $\lambda$ and $\mu$ characterize the hidden variables of the systems produced by the sources $\rho_{AB}$ and $\rho_{BC}$, respectively [75, 76]. Otherwise, it is called non-bilocal.
In order to detect non-bilocal correlations generated by the network, it is often necessary to find suitable measurements that violate the following bilocal inequality
(3.1) $\mathcal{S}\equiv\sqrt{|I|}+\sqrt{|J|}\leq 2,$
whose maximum quantum violation is $2\sqrt{2}$ and is attainable. Here
$I\equiv\sum_{x,z}\langle A_{x}B_{0}C_{z}\rangle=\langle(A_{0}+A_{1})B_{0}(C_{0}+C_{1})\rangle,$
$J\equiv\sum_{x,z}(-1)^{x+z}\langle A_{x}B_{1}C_{z}\rangle=\langle(A_{0}-A_{1})B_{1}(C_{0}-C_{1})\rangle$
as introduced in Ref. [77]:
$\langle A_{x}B_{y}C_{z}\rangle=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}{\rm tr}((A_{a|x}B_{b|y}C_{x|z})\rho_{AB}\otimes\rho_{BC})=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}p(abc|xyz).$
Here $A_{x}=\sum_{a}(-1)^{a}A_{a|x}$, $B_{y}=\sum_{b}(-1)^{b}B_{b|y}$ and $C_{z}=\sum_{z}(-1)^{z}C_{c|z}$, where $A_{a|x},\ B_{b|y}$, and $C_{c|z}$ are the positive operator-valued measurement (POVMs) performed by Alice, Bob and Charles, respectively.
Furthermore, it follows from Ref. [57] that the violation of the inequality
(3.2) $\mathcal{S}^{\prime}=|K|+|L|\leq 2$
can also be used to detect nonbilocal correlations, where $K=\langle A_{0}(B_{0}-B_{1})C_{0}\rangle$ and $L=\langle A_{1}(B_{0}+B_{1})C_{1}\rangle$ [78]. Notably, different inequalities exhibit varying capabilities in detecting nonlocality.
Mutually-commuting von Neumann algebra models. In QFT, the observables for Alice, Bob, and Charles are associated with three mutually-commuting von Neumann algebras $\mathcal{M}_{\mathcal{A}}$, $\mathcal{M}_{\mathcal{B}}$, $\mathcal{M}_{\mathcal{C}}$. Therefore, our model encompasses both the non-relativistic quantum mechanics scenario and the quantum field theory scenario. The idea of this model is similar to that in Refs. [61, 62].
Definition 1. (Ternary Mutually-commuting von Neumann Algebra Models of Tripartite Quantum Systems) Let $\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}}$ be von Neumann subalgebras of $\mathcal{B}(\mathcal{H})$ over some Hilbert space $\mathcal{H}$, which are mutually commuting, i.e., $\mathcal{M}_{i}\subset\mathcal{M}_{j}^{\prime}$ with $i\neq j\in\{\mathcal{A},\mathcal{B},\mathcal{C}\}$, where $\mathcal{M}_{j}^{\prime}$ is the communtant of $\mathcal{M}_{j}$. The generated von Neumann algebra
$\mathcal{M}_{\mathcal{ABC}}=(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}}\vee\mathcal{M}_{\mathcal{C}})^{\prime\prime}.$
We refer to the above model as the TMCvNA model. When $\mathcal{M}_{\mathcal{ABC}}\simeq\mathcal{M}_{\mathcal{A}}\otimes\mathcal{M}_{\mathcal{B}}\otimes\mathcal{M}_{\mathcal{C}}$, it is called the tensor product algebra model. In this paper, for any $A\in\mathcal{M}_{\mathcal{A}}$, $B\in\mathcal{M}_{\mathcal{B}}$, $C\in\mathcal{M}_{\mathcal{C}}$, we always assume that they are Hermitian.
We intend to use the above model of ternary mutually commuting von Neumann algebras to describe the entanglement swapping network in Fig. 1. We note that there is no correlation between the parties $\mathcal{A}$ and $\mathcal{C}$ in the network. Mathematically, this independence can be described by the following formula. So the network state of the entanglement swapping network $\tau$ should be a state in the dual space $\mathcal{M}_{\mathcal{ABC}}^{*}$, satisfying
() $\tau(AC)=\tau(A)\tau(C)$
for any $A\in\mathcal{M}_{\mathcal{A}},C\in\mathcal{M}_{\mathcal{C}}$. We call it the independent condition (*). This assumption will be used throughout this paper.
Definition 2. A ternary mutually-commuting von Neumann algebra model of entanglement swapping networks is the ternary mutually-commuting von Neumann algebra model of tripartite quantum systems with the independence condition (*).
4. Bilocal inequalities and their bounds
In this section, we further analyze the conditions under which the bilocal inequality holds or is violated in the TMCvNA model of an entanglement swapping network. Specifically, in this model, we can construct bilocal inequalities analogous to those in the non-relativistic setting. Define
$I_{\tau}=\tau\bigl((A_{0}+A_{1})B_{0}(C_{0}+C_{1})\bigr),\ J_{\tau}=\tau\bigl((A_{0}-A_{1})B_{1}(C_{0}-C_{1})\bigr),$
where $\tau$ is the state on $\mathcal{M}_{\mathcal{ABC}}$ satisfying the independent condition (*). Here, $\tau\bigl(A_{x}B_{y}C_{z}\bigr)=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}\tau({A}_{a|x}{B}_{b|y}{C}_{c|z})$ and $A_{x}=\sum_{a}(-1)^{a}A_{a|x}$, $B_{y}=\sum_{b}(-1)^{b}B_{b|y}$, $C_{z}=\sum_{z}(-1)^{z}C_{c|z}$, where $A_{a|x}$, $B_{b|y}$, and $C_{c|z}$ are the POVMs performed by Alice, Bob, and Charles, respectively. Moreover, the network correlation $\hat{p}=p(\alpha\beta\gamma|xyz)$ in the TMCvNA model is defined as
$p(\alpha\beta\gamma|xyz)=\tau({A}_{\alpha|x}{B}_{\beta|y}{C}_{\gamma|z}),$
In the TMCvNA model, analogous to Ineq. (3.1), we set
(4.1) $\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}.$
Analogously, corresponding to Ineq. (3.2), we define
(4.2) $\mathcal{S}_{\tau}^{\prime}=|K_{\tau}|+|L_{\tau}|,$
where, $K_{\tau}=\tau(A_{0}(B_{0}+B_{1})C_{0}),\ L_{\tau}=\tau(A_{1}(B_{0}-B_{1})C_{1})$.
We say that the state $\tau$ together with the observables $A_{x}$, $B_{y}$, $C_{z}$ satisfies the bilocal inequality if $\mathcal{S}_{\tau}\leq 2$, and violates it if $2<\mathcal{S}_{\tau}\leq 2\sqrt{2}$. The same criterion applies to $\mathcal{S}_{\tau}^{\prime}$ in Ineq. (4.2).
Next, we rely on the abelianness of the algebra to determine the bounds for the bilocal inequality, respectively.
The following conclusion indicates that in the mutually-commuting von Neumann algebra model, the supremum of $\mathcal{S}_{\tau}$ defined in Ineq. (4.1) is $2\sqrt{2}$. This coincides with the case in non-relativistic quantum mechanics, where the bilocal quantity $\mathcal{S}$ in Ineq. (3.1) attains a maximal violation of $2\sqrt{2}$ based on quantum resources.
Theorem 1. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, $\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2\sqrt{2}.$
Proof. According to the Gelfand-Namark-Segal (GNS) construction, there is a -representation $\pi_{\tau}:\mathcal{M}_{\mathcal{ABC}}\rightarrow\mathcal{B}(\mathcal{H}_{\tau})$ and a cyclic vector $\Omega\in\mathcal{B}(\mathcal{H}_{\tau})$ such that the set $\{\pi_{\tau}(O)\Omega:\ O\in\mathcal{M}_{\mathcal{ABC}}\}$ is dense in $\mathcal{H}_{\tau}$. It follows by applying the Cauchy-Schwarz inequality that
$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}$
$=\sqrt{|\tau((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\tau((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}$
$\leq\sqrt{2}\sqrt{|\tau(B_{0}(A_{0}+A_{1})(C_{0}+C_{1}))|+|\tau(B_{1}(A_{0}-A_{1})(C_{0}-C_{1}))|}$
$=\sqrt{2}\sqrt{|\langle\pi_{\tau}(B_{0})\Omega,\pi_{\tau}((A_{0}+A_{1})(C_{0}+C_{1}))\Omega\rangle|+|\langle\pi_{\tau}(B_{1})\Omega,\pi_{\tau}((A_{0}-A_{1})(C_{0}-C_{1}))\Omega\rangle|}$
$\leq\sqrt{2}\sqrt{\|\pi_{\tau}(B_{0})\Omega\|\|\pi_{\tau}((A_{0}+A_{1})(C_{0}+C_{1}))\Omega\|+\|\pi_{\tau}(B_{1})\Omega\|\|\pi_{\tau}((A_{0}-A_{1})(C_{0}-C_{1}))\Omega\|}$
$\leq\sqrt{2}\sqrt{\sqrt{\tau((A_{0}+A_{1})^{2}(C_{0}+C_{1})^{2})}+\sqrt{\tau((A_{0}-A_{1})^{2}(C_{0}-C_{1})^{2})}}$
$\leq\sqrt{2}\sqrt{\sqrt{\tau(A_{0}+A_{1})^{2}+\tau(A_{0}-A_{1})^{2}}\sqrt{\tau(C_{0}+C_{1})^{2}+\tau(C_{0}-C_{1})^{2}}}$
$=\sqrt{2}\sqrt{\sqrt{2\tau(A_{0}^{2}+A_{1}^{2})}\sqrt{2\tau(C_{0}^{2}+C_{1}^{2})}}$
$\leq\sqrt{2}\sqrt{2\sqrt{2}\sqrt{2}}$
$=2\sqrt{2}.$
[…]
Building on the results above, we now investigate how the quantity $\mathcal{S}_{\tau}$ in Ineq. (4.1) depends on the abelianness of the algebras $\mathcal{M}_{\mathcal{A}}$, $\mathcal{M}_{\mathcal{B}}$, and $\mathcal{M}_{\mathcal{C}}$. Specifically, the results indicate that in entanglement swapping networks, the abelianness of the three algebras plays distinct roles in reducing the upper bound of the inequality to 2, i.e., determining the conditions under which no violation of the bilocal inequality can occur. This is not a simple generalization of the bipartite Bell scenario [27], where, with only two systems, Summers et al. showed that if one of these two algebras is abelian, the upper bound of the Bell inequality is 2.
Theorem 2. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks,
(1) if $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, then
$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2;$
(2) if $\mathcal{M}_{\mathcal{B}}$ is Abelian, then
$\mathcal{S}_{\tau}^{\prime}=|K_{\tau}|+|L_{\tau}|\leq 2.$
Proof. To see (1), since $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, respectively, the eight elements
$A_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{0}A_{0})(1+\epsilon_{1}A_{1}),\ \ \ C_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{1}C_{0})(1+\epsilon_{1}C_{1})$
with $\epsilon_{0},\epsilon_{1}\in\{+,-\}$ are positive. By direct computation, one obtains that
$A_{0}+A_{1}=2(A_{++}-A_{--}),\ C_{0}+C_{1}=2(C_{++}-C_{--}),$
$A_{0}-A_{1}=2(A_{+-}-A_{-+}),\ C_{0}-C_{1}=2(C_{+-}-C_{-+}).$
So one obtains that
$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\tau((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}$
$=2\left(\sqrt{|\tau((A_{++}-A_{--})B_{0}(C_{++}-C_{--}))|}+\sqrt{|\tau((A_{+-}-A_{-+})B_{1}(C_{+-}-C_{-+}))|}\right)$
$=2\left(\sqrt{|\tau(A_{++}B_{0}C_{++})-\tau(A_{++}B_{0}C_{--})-\tau(A_{--}B_{0}C_{++})+\tau(A_{--}B_{0}C_{--})|}\right$.
$\qquad\left.+\sqrt{|\tau(A_{+-}B_{1}C_{+-})-\tau(A_{+-}B_{1}C_{-+})-\tau(A_{-+}B_{1}C_{+-})+\tau(A_{-+}B_{1}C_{-+})|}\right)$
$\leq 2\left(\sqrt{|\tau(A_{++}C_{++})+\tau(A_{++}C_{--})+\tau(A_{--}C_{++})+\tau(A_{--}C_{--})|}\right$.
$\qquad\left.+\sqrt{|\tau(A_{+-}C_{+-})+\tau(A_{+-}C_{-+})+\tau(A_{-+}C_{+-})+\tau(A_{-+}C_{-+})|}\right)$
$=2\left(\sqrt{|\tau(A_{++}+A_{--})||\tau(C_{++}+C_{--})|}+\sqrt{|\tau(A_{+-}+A_{-+})||\tau(C_{+-}+C_{-+})|}\right)$
$\leq 2\sqrt{\tau(A_{++}+A_{--})+\tau(A_{+-}+A_{-+})}\sqrt{\tau(C_{++}+C_{--})+\tau(C_{+-}C_{-+})}$
$=2\sqrt{1}\sqrt{1}=2,$
where the first inequality follows the fact that $I\leq B_{i}\leq I\ (i=0,1)$ and the order-preserving of state $\tau$, and the second inequality holds because of the Cauchy Schwarz inequality and the non-negativeness of $A_{\epsilon_{0}\epsilon_{1}}\ \text{and}\ \ B_{\epsilon_{0}\epsilon_{1}}$.
To prove (2), since $\mathcal{M}_{\mathcal{B}}$ is Abelian, the four elements
$B_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{0}a_{0})(1+\epsilon_{1}a_{1}),$
with $\epsilon_{0},\epsilon_{1}\in\{+,-\}$ are positive. It follows from a direct computation that
$B_{0}+B_{1}=2(B_{++}-B_{--}),\ B_{0}-B_{1}=2(B_{+-}-B_{-+}).$
So one obtains that
$\mathcal{S}_{\tau}^{\prime}$ $=$ $\tau[A_{0}(B_{0}+B_{1})C_{0}]+\tau[A_{1}(B_{0}-B_{1})C_{1}]$
$=$ $2(\tau[A_{0}(B_{++}-B_{--})C_{0}]+\tau[A_{1}(B_{+-}-B_{-+})C_{1}])$
$\leq$ $2(\tau[B_{++}]+\tau[B_{--}]+\tau[B_{+-}]+\tau[B_{-+}])=2$
[…]
Remark. The above theorem illustrates a phenomenon: the violations of the two Bell-like inequalities, i.e., $2<\mathcal{S}_{\tau}\leq 2\sqrt{2}$ and $2<\mathcal{S}^{\prime}_{\tau}\leq 2\sqrt{2}$ can serve as an indicator of the non-abelianness of the underlying algebras. Applying the theorem, we can infer from a violation of Ineq. (4.1) that at least one of the algebras $\mathcal{M_{A}}$ and $\mathcal{M_{C}}$ is non-abelian; Similarly, a violation of Ineq. (4.2) implies that $\mathcal{M_{B}}$ is non-abelian. Here, one also needs to consider the chosen observables and the specific state.
Nevertheless, it is important to note that these conclusions depend not only on the algebraic structure but also on the specific choice of observables and the particular state under consideration. To isolate the influence of the algebraic structure and examine separately the role of observables and the state in determining the values $\mathcal{S}_{\tau}$ and $\mathcal{S}_{\tau}^{\prime}$ in Ineqs. (4.1) and  (4.2), the following conclusion can be regarded as a generalization of the results in the non-relativistic quantum mechanical framework. For the sake of clarity, we introduce a tensor product structure into the TMCvNA model. Before presenting the following result, we first introduce the definition of separable states.
Definition 3. (Separable states) Let $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{B}}$ be mutually commuting von Neumann subalgebras of $\mathcal{M}_{\mathcal{AB}}=(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}})^{\prime\prime}$. A state $\tau\in\mathcal{M}_{\mathcal{AB}}^{*}$ is said to be separable if there exist states $\xi_{\alpha}\in\mathcal{M}_{\mathcal{A}}^{*}$, $\eta_{\alpha}\in\mathcal{M}_{\mathcal{B}}^{*}$, and positive coefficients $\lambda_{\alpha}$ with $\sum_{\alpha}\lambda_{\alpha}=1$ such that for all $A\in\mathcal{M}_{\mathcal{A}},B\in\mathcal{M}_{\mathcal{B}}$,
$\tau(AB)=\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(A)\eta_{\alpha}(B).$
[…]
$\sup_{\{A_{x},B_{y},C_{z}\}}\mathcal{S}_{\tau}=\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|})=2.$
[…]
$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\tau((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}$
$=\sqrt{|\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(A_{0}+A_{1})\eta_{\alpha}(B_{0}^{L})||\sum_{\beta}\mu_{\beta}M_{\beta}(B_{0}^{R})N_{\beta}(C_{0}+C_{1})|}$
$\hskip 18.49988pt+\sqrt{|\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(A_{0}-A_{1})\eta_{\alpha}(B_{1}^{R})||\sum_{\beta}\mu_{\beta}M_{\beta}(B_{1}^{R})N_{\beta}(C_{0}-C_{1})|}$
$\leq\sqrt{(\sum_{\alpha}\lambda_{\alpha}|\xi_{\alpha}(A_{0}+A_{1})|)(\sum_{\beta}\mu_{\beta}|N_{\beta}(C_{0}+C_{1})|)}+\sqrt{(\sum_{\alpha}\lambda_{\alpha}|\xi_{\alpha}(A_{0}-A_{1})|)(\sum_{\beta}\mu_{\beta}|N_{\beta}(C_{0}-C_{1})|)}$
$\leq\sqrt{\sum_{\alpha}\lambda_{\alpha}(|\xi_{\alpha}(A_{0}+A_{1})|+|\xi_{\alpha}(A_{0}-A_{1})|)}\sqrt{\sum_{\beta}\mu_{\beta}(|N_{\beta}(C_{0}+C_{1})|+|N_{\beta}(C_{0}-C_{1})|)}$
$=2\sqrt{\sum_{\alpha}\lambda_{\alpha}\max\{\xi_{\alpha}(A_{0}),\xi_{\alpha}(A_{1})\}}\sqrt{\sum_{\beta}\mu_{\beta}\max\{N_{\beta}(C_{0}),N_{\beta}(C_{1})\}}$
$\leq 2.$
[…]
From the above conclusion, we see that whether a bilocal inequality can be violated depends on the choice of observables and the state, and even more so on the structure of the algebra. In the following, we define a quantity
$\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|})$
that appears in Theorem 3.
Combining Theorems 1, 2 and 3, one naturally obtains the following corollary. A similar conclusion can be drawn for the quantity $\mathcal{S}^{\prime}_{\tau}$ in Ineq. (4.2).
Corollary 4. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks,
(1) for any state $\tau$ and any choice of observables $A_{x}\in\mathcal{M}_{\mathcal{A}},\ B_{y}\in\mathcal{M}_{\mathcal{B}},\ C_{z}\in\mathcal{M}_{\mathcal{C}}$ in a scheme with two inputs and two outputs, the quantity $\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})$ satisfies
$2\leq\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})\leq 2\sqrt{2}.$
(2) if $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, then $\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=2$.
(3) for any states $\phi,\psi\in[(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}}\vee\mathcal{M}_{\mathcal{C}})^{\prime\prime}]^{*}$, the following inequality holds:
$|\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})-\mathcal{S}(\psi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})|\leq k\sqrt{\|\phi-\psi\|},$
where $k$ is a positive constant. Consequently, the functional $\phi\rightarrow\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|})$ is norm continuous.
Proof. To show (1), note that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and combining this with the proof of Theorem 2, we obtain (1).
(2) holds by the fact that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and by Theorem 2.
To prove (3), note that by the representations of $I_{\tau}$ and $J_{\tau}$, together with the facts that $|\sup x-\sup y|\leq\sup|x-y|$ and the triangle inequality, one can obtain
$|\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})-\mathcal{S}(\psi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})|$
$=\left|\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|}\right)-\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{|I_{\psi}|}+\sqrt{|J_{\psi}|}\right)\right|$
$=\left|\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{|\phi((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\phi((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}\right)\right$.
$\qquad\left.-\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{|\psi((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\psi((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}\right)\right|$
$\leq\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left|\sqrt{|\phi((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\phi((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}\right$.
$\qquad\left.-\sqrt{|\psi((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}-\sqrt{|\psi((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}\right|$
$\leq\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\left|\sqrt{|\phi((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}-\sqrt{|\psi((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}\right|\right$.
$\qquad\left.+\left|\sqrt{|\phi((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}-\sqrt{|\psi((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}\right|\right)$
$\leq\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{|(\phi-\psi)((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|(\phi-\psi)((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}\right)$
$\leq\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{\|\phi-\psi\|\,\|(A_{0}+A_{1})B_{0}(C_{0}+C_{1})\|}+\sqrt{\|\phi-\psi\|\,\|(A_{0}-A_{1})B_{1}(C_{0}-C_{1})\|}\right)$
$\leq k\sqrt{\|\phi-\psi\|},$
[…]
5. Maximal violation of bilocal inequalities and algebraic structures
In this section, we aim to point out that the violation of bilocal inequalities, in particular the maximal violation, can reflect the structural properties of the algebra. Here, violation refers to exceeding the upper bound of 2 for the $\mathcal{S}_{\tau}$ in inequality (4.1), while maximal violation means attaining the value $2\sqrt{2}$ for the same quantity.
In the following theorem, we attempt to analyze the conditions for maximal violation.
Theorem 5. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if $\tau$ is faithful, then
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$
if and only if $\tau(A_{i}^{2}A)=\tau(A),\ \tau(B_{i}^{2}B)=\tau(B),\ \tau(C_{i}^{2}C)=\tau(C)$, and $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=0,\ \tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$ for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$ with $i\in\{0,1\}$.
Remark. This theorem states that when the inequality is maximally violated, $A_{0},\ A_{1}$ and $\frac{i}{2}[A_{0},A_{1}]$ form a realization of the Pauli spin matrices in $\mathcal{M}_{\mathcal{A}}$ and generate a copy of $M_{2}(\mathbb{C})$ in $\mathcal{M}_{\mathcal{A}}$. The same holds for $C_{0},\ C_{1}$ and $\frac{i}{2}[C_{0},C_{1}]$ in $\mathcal{M}_{\mathcal{C}}$. Moreover, $\tau(A_{0}A_{1}+A_{1}A_{0})=0,\ \tau(C_{0}C_{1}+C_{1}C_{0})=0.$
Proof. If $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$, it follows from the proof of Theorem 1 that for any $t\in[0,1]$, these equalities hold:
(5.1) $|\tau(B_{0}(A_{0}+A_{1})(C_{0}+C_{1}))|=|\tau(B_{1}(A_{0}-A_{1})(C_{0}-C_{1}))|,$
(5.2) $\pi_{\tau}(B_{0})\Omega=k_{0}\pi_{\tau}[(A_{0}+A_{1})(C_{0}+C_{1})]\Omega,$
(5.3) $\pi_{\tau}(B_{1})\Omega=k_{1}\pi_{\tau}[(A_{0}-A_{1})(C_{0}-C_{1})]\Omega,$
(5.4) $\|\pi_{\tau}(B_{0})\Omega\|=\|\pi_{\tau}(B_{1})\Omega\|=1,$
(5.5) $\tau[(A_{0}+A_{1})^{2}]=t\tau[(C_{0}+C_{1})^{2}],$
(5.6) $\tau[(A_{0}-A_{1})^{2}]=t\tau[(C_{0}-C_{1})^{2}],$
(5.7) $A_{0}^{2}+A_{1}^{2}=2I,C_{0}^{2}+C_{1}^{2}=2I.$
From (5.7), one gets $A_{i}^{2}=I,\ C_{i}^{2}=I\ (i=0,1)$ because $I\leq A_{i},C_{i}\leq I$, therefore $\tau(A_{i}^{2}A)=\tau(A),\ \tau(C_{i}^{2}C)=\tau(C)$ for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$.
Now, let us show the proof for $\tau(B_{i}^{2}B)=\tau(B),\ \tau[(A_{0}A_{1}+A_{1}A_{0})A]=0$, and $\tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$ for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$, which implies $B_{i}^{2}=I,$ $A_{0}A_{1}+A_{1}A_{0}=0,$ and $C_{0}C_{1}+C_{1}C_{0}=0$. According to Eqs. (5.5)-(5.7), i.e.,
$\left\{\begin{array}[]{c}\tau(A_{0}^{2}+A_{1}^{2}+A_{0}A_{1}+A_{1}A_{0})=t\tau(C_{0}^{2}+C_{1}^{2}+C_{0}C_{1}+C_{1}C_{0})\\ \tau(A_{0}^{2}+A_{1}^{2}-A_{0}A_{1}-A_{1}A_{0})=t\tau(C_{0}^{2}+C_{1}^{2}-C_{0}C_{1}-C_{1}C_{0})\\ A_{0}^{2}+A_{1}^{2}=2I,\ C_{0}^{2}+C_{1}^{2}=2I,\end{array}\right$.
we can get
$t=1,\ \tau(A_{0}A_{1}+A_{1}A_{0})=\tau(C_{0}C_{1}+C_{1}C_{0}).$
Then combing conditions (5.2), (5.3) and condition (5.4), one can get
$k_{0}^{2}\tau[(A_{0}+A_{1})^{2}(C_{0}+C_{1})^{2}]=1,\ k_{1}^{2}\tau[(A_{0}-A_{1})^{2}(C_{0}-C_{1})^{2}]=1,$
i.e.,
(5.8) $|k_{0}|=\frac{1}{2+\tau(X)},\ \ \ |k_{1}|=\frac{1}{2-\tau(X)}$
because of $\tau(X)=\tau(Y)$, where $X=A_{0}A_{1}+A_{1}A_{0},\ Y=C_{0}C_{1}+C_{1}C_{0}$. Then according to Eqs. (5.2), (5.3), and (5.1),
$|\tau[B_{0}(A_{0}+A_{1})(C_{0}+C_{1})]|=|k_{0}||\tau[(A_{0}+A_{1})^{2}(C_{0}+C_{1})^{2}]|,$
$|\tau[B_{1}(A_{0}-A_{1})(C_{0}-C_{1})]|=|k_{1}||\tau[(A_{0}-A_{1})^{2}(C_{0}-C_{1})^{2}]|.$
Substituting (5.8) and (5.1) to the above equations, we have
$\frac{1}{2+\tau(X)}(2+\tau(X))^{2}=\frac{1}{2-\tau(X)}(2-\tau(X))^{2}$
deriving $\tau(X)=\tau(Y)=0$. It follows from $\tau(X)=0=\tau(Y)$ and $\|\pi_{\tau}(B_{i})\Omega\|=1$ that
$k_{0}^{2}=k_{1}^{2}=\frac{1}{4}$
from Eq. (5.8). Since
$|\langle\Omega,\pi_{\tau}(B_{0}^{2})\Omega\rangle|$ $=$ $|\frac{1}{4}\langle\Omega,\pi_{\tau}[(A_{0}+A_{1})^{2}(C_{0}+C_{1})^{2}]\Omega\rangle|$
$\leq$ $\frac{1}{4}\sqrt{\|\pi_{\tau}[(A_{0}+A_{1})(C_{0}+C_{1})]\Omega\|}\sqrt{\|\pi_{\tau}[(A_{0}+A_{1})(C_{0}+C_{1})]\Omega\|}$
$=$ $\frac{1}{4}\tau[(A_{0}+A_{1})^{2}]\tau[(C_{0}+C_{1})^{2}]=1,$
and $\tau(B_{0}^{2})=|\langle\Omega,\pi_{\tau}(B_{0}^{2})\Omega\rangle|=1$, this implies that $\pi_{\tau}[(B_{0}^{2})]\Omega=\Omega$, so
$\tau(B_{0}^{2}B)=\langle\Omega,\pi_{\tau}(B_{0}^{2})\pi_{\tau}(B)\Omega\rangle=\langle\pi_{\tau}(B_{0}^{2})\Omega,\pi_{\tau}(B)\Omega\rangle=\langle\Omega,\pi_{\tau}(B)\Omega\rangle=\tau(B).$
Similarly $\tau(B_{1}^{2}B)=\tau(B)$ for any $B\in\mathcal{M}_{\mathcal{B}}.$ Furthermore, it follows from Eq. (5.2) and $\pi_{\tau}[(B_{0}^{2})]\Omega=\Omega$ that
$\pi_{\tau}(2X+2Y+XY)\Omega=0.$
So $\tau[(2X+2Y+XY)^{2}]=0,$ implying that
$\tau(X^{2})=\tau(Y^{2})=0.$
Combining the faithfulness, non-negativity of the state $\tau$ and the self-adjointness of X,YX,\ Y. Then $X=Y=0$, i.e.,
$A_{0}A_{1}+A_{1}A_{0}=C_{0}C_{1}+C_{1}C_{0}=0,$
and implies that $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=\tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$ .
[…]
The above Theorem 5 indicates that if the bilocal inequality is maximally violated, then the underlying algebra locally contains an $M_{2}(\mathbb{C})$ factor. In contrast, the follwing theorem presents a highly specialized scenario in which a precise characterization of algebras attaining maximal viloation can be obtained. Note that next we denote by $\Lambda_{1}\overline{\otimes}\Lambda_{2}$ the von Neumann algebra of the tensor product of $\Lambda_{1},\Lambda_{2}$
Theorem 6. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if there are hyperfinite type II1 factors $\mathcal{R}_{\mathcal{A}}$ and $\mathcal{R}_{\mathcal{C}}$ such that $\mathcal{M}_{\mathcal{A}}\simeq\mathcal{M}_{\mathcal{A}}\overline{\otimes}\mathcal{R}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}\simeq\mathcal{M}_{\mathcal{C}}\overline{\otimes}\mathcal{R}_{\mathcal{C}}$, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$.
Proof. We first deal with the special case: $\mathcal{M}_{\mathcal{A}}=\mathcal{R}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}=\mathcal{R}_{\mathcal{C}}$. Let $\widetilde{\mathcal{H}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{H}_{\alpha},\Omega_{\alpha})$ be the incomplete tensor product of $\mathcal{H}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Omega_{\alpha}=\frac{1}{\sqrt{2}}(\Phi_{1}\otimes\Phi_{1}+\Phi_{2}\otimes\Phi_{2})\equiv\Omega\text{ for some basis }\{\Phi_{1},\Phi_{2}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then by the construction of Araki and Woods [79], $\mathcal{R}_{\mathcal{A}}\simeq\mathcal{R}(\mathcal{H}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Omega_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{H}}$. Similarly, let $\widetilde{\mathcal{K}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{K}_{\alpha},\Lambda_{\alpha})$ be the incomplete tensor product of $\mathcal{K}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Lambda_{\alpha}=\frac{1}{\sqrt{2}}(\Psi_{3}\otimes\Psi_{3}+\Psi_{4}\otimes\Psi_{4})\equiv\Lambda\text{ for some basis }\{\Psi_{3},\Psi_{4}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then $\mathcal{R}_{\mathcal{C}}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{K}}$. We can obtain that there is $\mathcal{R}_{\mathcal{C}}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$. Now, from Theorem 5, we know that $A_{0},\ A_{1}$ and $A_{2}=-\frac{i}{2}[A_{0},A_{1}]$ form a realization of the Pauli spin matrices on $\mathcal{H}_{\alpha}$, and the same holds for $C_{0},\ C_{1}$ and $C_{2}=-\frac{i}{2}[C_{0},C_{1}]$ in $\mathcal{K}_{\alpha}$, and $B_{0},B_{1}$ with $B_{i}^{2}=I$ such that the quantity $\mathcal{S}_{\tau}$ in Ineq. (4.1) arrives at $2\sqrt{2}$. Applying similar technologies of Eqs. (6)-(8) of Ref.[77], one can obtain that
(5.9) $\sqrt{|\langle\Omega\otimes\Lambda,[(A_{0}+A_{1})B_{0}(C_{0}+C_{1})]\Omega\otimes\Lambda\rangle|}=\sqrt{2},$
(5.10) $\sqrt{|\langle\Omega\otimes\Lambda,[(A_{0}-A_{1})B_{1}(C_{0}-C_{1})]\Omega\otimes\Lambda\rangle|}=\sqrt{2}.$
[…]
$\widetilde{T}_{\alpha}=(\widetilde{A}_{0,\alpha}+\widetilde{A}_{1,\alpha})\widetilde{B}_{0,\alpha}(\widetilde{C}_{0,\alpha}+\widetilde{C}_{1,\alpha}),$
$\widetilde{S}_{\alpha}=(\widetilde{A}_{0,\alpha}-\widetilde{A}_{1,\alpha})\widetilde{B}_{1,\alpha}(\widetilde{C}_{0,\alpha}-\widetilde{C}_{1,\alpha}).$
Furthermore, $\Xi=\bigotimes_{\alpha\in\mathbb{N}}\Xi_{\alpha},\Upsilon=\bigotimes_{\alpha\in\mathbb{N}}\Upsilon_{\alpha},$ with $\Xi_{\alpha}=\Omega_{\alpha}\otimes\Omega_{\alpha}$ and $\Upsilon_{\alpha}=\Lambda_{\alpha}\otimes\Lambda_{\alpha}$ for almost all $\alpha$. By the definition of the incomplete infinite tensor product, such vectors are total in $\widetilde{\mathcal{H}}\otimes\widetilde{\mathcal{K}}$. But one has
$\sqrt{\langle\Xi,\widetilde{T}_{\alpha}\Upsilon\rangle}=\langle\Xi,\Upsilon\rangle\cdot\sqrt{2},\sqrt{\langle\Xi,\widetilde{S}_{\alpha}\Upsilon\rangle}=\langle\Xi,\Upsilon\rangle\cdot\sqrt{2}.$
for all $\alpha$ larger than some $\alpha_{0}$ (depending on $\Xi$ and $\Upsilon$).
[…]
It is mentioned that the results of Theorem 6 can be applied to quantum field theory. In the algebraic framework of the QFT, wedge algebras (von Neumann algebras associated to specific unbounded regions in Minkowski spacetime, such as $\mathcal{O}$ associated to tangent regions, are typically type III1 factors. By the classical result of A. Connes, all injective infinite factors (except some type III0) are strongly stable (meaning they can absorb a hyperfinite II1 factor: $\mathcal{M}\simeq\mathcal{M}\overline{\otimes}\mathcal{R}_{1}$).
Corollary 7. Let $\mathcal{O}_{i}$ $(i=1,2,3)$ be wedge-shaped regions, $\mathcal{A}(\mathcal{O}_{i})$ be the wedge algebra on a separable Hilbert space $H$ with cyclic and separating vector. If $\mathcal{A}(\mathcal{O}_{1})$ and $\mathcal{A}(\mathcal{O}_{3})$ are both type III1 factors, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$ on $(\mathcal{A}(\mathcal{O}_{1})\vee\mathcal{A}(\mathcal{O}_{2})\vee\mathcal{A}(\mathcal{O}_{3}))^{\prime\prime}$.
Acknowledgement. Authors thank for your comments. This work is supported by the National Natural Science Foundation of China (Grant No. 12271394).