### v5

# B. Zheng, S. Yang, J. Hou and K. He

[…]
B. Zheng, S. Yang, J. Hou and K. He
[…]
Violation of Bell-type Inequalities on Mutually-commuting von Neumann Algebra Models
College of Mathematics, Taiyuan University of Technology, Taiyuan, 030024, China
School of Mathematics, North University of China, Taiyuan, 030051, China,
School of Cyberspace Science and Technology, Beijing Institute of Technology, Beijing, 100081, China
E-mail :: yangshuyuan2000@163.com
College of Mathematics, Taiyuan University of Technology, Taiyuan, 030024, China
E-mail :: jinchuanhou@aliyun.com
College of Mathematics, Taiyuan University of Technology, Taiyuan, 030024, China
E-mail :: hekan@tyut.edu.cn
Violation of Bell-type Inequalities on Mutually-commuting von Neumann Algebra Models of Entanglement Swapping Networks  [^1] Shuyuan Yang and Bingke Zheng contribute equally to the work. Supported by the National Natural Science Foundation of China (Grant No. 12271394, 12571138)
Bingke ZHENG    Shuyuan YANG    Jinchuan HOU    Kan HE
[…]
Violation of Bell inequalities in bipartite systems represented by mutually-commuting von Neumann algebras has pioneered the study of vacuum entanglement in algebraic quantum field theory. It is unexpected that the maximal violation of Bell inequality can discover algebraic structures. In the paper, we establish the mutually-commuting von Neumann algebra model for entanglement swapping networks and Bell-type inequalities on this model. It generalizes the bipartite case to the ternary case. These algebras are all general von Neumann algebras, which provide a natural perspective to investigate Bell nonlocality in quantum networks in the infinitely-many-degree-of-freedom setting. We determine various bounds for Bell-type inequalities based on the structure of von Neumann algebras, and identify the algebraic structural conditions required for their violation. Finally, we show that the maximal violation of Bell-type inequalities in entanglement swapping networks can be used to determine partially the type classification of the underlying von Neumann algebras.
[…]
von Neumann algebra, Bell nonlocality, Bell inequality, Quantum network
[…]
81R15, 47A63, 47L30, 47N50
1 Introduction
In non-relativistic quantum mechanics, Bell nonlocality demonstrates that local measurements performed on one subsystem of a quantum state can instantaneously influence the measurement outcomes on another subsystem, regardless of the spatial separation between them [1, 2, 3]. Such nonlocal correlations can be detected through Bell inequalities, which serve as constraints that all local correlations must obey [4, 5, 6, 7, 8]. It has been demonstrated to offer quantum advantages in various device-independent quantum information tasks, including communication complexity [9], quantum key distribution [10, 11], randomness amplification [12, 13], and measurement-based quantum computation [14, 15].
Meanwhile, motivated by the quantum field theory (QFT), which originates from the study of relativistic quantum mechanics, many novel quantum phenomena in systems with infinitely many degrees of freedom have been discovered [16, 17, 18, 19, 20, 21, 22]. This differs from the non-relativistic quantum-mechanical setup, which is usually linked to type I von Neumann algebras and relies on the algebraic tensor product as its mathematical framework [23, 24, 25]. These are two distinct models, referred to respectively as the tensor product algebra (TPA) model and the mutually-commuting von Neumann (observable) algebra (MCvNA) model. In the MCvNA model, there is, in general, no tensor product decomposition of the Hilbert space describing subsystems. However, it should be pointed out that relying solely on the TPA model to discuss quantum information problems has drawbacks [26, 27, 28]. It fails to provide a universal framework for accurately describing phenomena in systems with infinite degrees of freedom and the quantum field theory, which requires the language of type III von Neumann algebras. Research on quantum information problems on von Neumann algebras has received significant attention and yielded many meaningful results from a mathematical perspective [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43].
In the MCvNA model, the algebra of observables of quantum systems is described by a von Neumann algebra $\mathcal{M}$, with $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{B}}$ being two mutually commuting von Neumann subalgebras of $\mathcal{M}$ such that $(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}})^{\prime\prime}=\mathcal{M}$. Here, $\mathcal{M}^{\prime\prime}$ denotes the double commutant of $\mathcal{M}$ [44, 45, 19]. It has been shown that the mutually-commuting von Neumann algebra model provides a more general framework [19]. In the 1980s early, Summers et al. first introduced the maximal violation of Bell inequality and proved that its value is bounded by $2\sqrt${2} in the MCvNA model of bipartite systems, with equality attainable if and only if each algebra contains a copy of $\mathcal{M}_{2}({\mathbb{C}})$ [46]. This shows that Bell nonlocality is not merely a quantum peculiarity but a structural feature encoded in the classification of operator algebras, providing rigorous tools to quantify non-classical correlations in relativistic quantum systems [47]. Translating these bounds into the vacuum representation of algebraic quantum field theory, they show that tangent wedge algebras are always maximally correlated, whereas strictly spacelike-separated wedges decay exponentially with mass-governed distance [48, 49, 50, 51, 52]. These works reveal a novel algebraic invariant, termed the Bell correlation invariant, which distinguishes infinitely many isomorphism classes of pairs of mutually commuting von Neumann algebras and links the maximal violation to the occurrence of the hyperfinite type $\rm II_{1}$ factor [47]. This is a pioneering work to make Bell nonlocality in QFT serve as a crucial bridge connecting quantum information science with fundamental physics [47, 53]. It provides a rigorous framework for reconciling quantum entanglement with relativistic causality, resolves conceptual challenges such as impossible measurements, and reveals how fundamental symmetries like parity violation affect quantum correlations [54, 55, 56, 57, 58].
In contrast to entanglement originating from an individual source, quantum networks comprise numerous small-scale entangled states. Owing to the independence among distinct sources, the correlations emerging from quantum networks exhibit non-convex characteristics that transcend the polytopes associated with single-source entanglement [59, 60, 61, 62, 63, 64]. To date, Bell-type inequalities in the non-relativistic quantum mechanics have been devised to certify nonlocal correlations across diverse network architectures, such as entanglement-swapping networks [59, 65], chain configurations [66, 67], star topologies [68, 69], polygon structures [70, 71, 72], tree-shaped networks [73, 74, 75], arbitrary acyclic networks [61, 62, 76], and arbitrary $k$-independent networks [77]. Alternative research directions examine the stronger forms of network nonlocality that surpass hybrid implementations involving classical variables and post-quantum resources [78, 79]. Nevertheless, limited progress has been made concerning the discrimination of correlations produced by different networks and the subsequent identification of underlying quantum network topologies [80]. Recently, the notion of bilocality in an entanglement swapping network based on the MCvNA model has already been introduced by Ligthart et al. [81, 82], and Xu has addressed the inclusion problem between TPA model and MCvNA model in this setting [83]. However, Bell-type inequalities in the MCvNA model have not yet been established. In this paper, we aim to establish bilocal inequalities within the mutually-commuting von Neumann algebra model and investigate how the degree of their violation is related to the structural properties of the algebras.
2 Ternary mutually-commuting von Neumann algebra models and entanglement swapping networks
Figure 1: An entanglement swapping network scenario with two sources. The connection between two parties represents the sharing of the physical system between them.
Quantum bilocal scenario. In non-relativistic quantum mechanics, the quantum entanglement swapping network (see Fig. 1) is a scenario of three parties consisting of Alice, Bob and Charles, and two sources $\rho_{AB}$, $\rho_{BC}$ shared between them. The inputs and outputs of the measurements performed by the three parties are denoted as $x,y,z$ and $a,b,c$, respectively. Assume that each party performs binary-input and binary-output measurements, with the observables for Alice, Bob, and Charles denoted as $A_{x}$, $B_{y}$ and $C_{z}$, respectively, with $x,y,z,a,b,c\in\{0,1\}$. Here it is required that the spectra of operators $A_{x},B_{y},C_{z}$ are all $\{-1,1\}$, implying that $I\leq A_{x},B_{y},C_{z}\leq I$. The correlations between the measurement outcomes of the three parties are described by the joint probability distribution $p(abc|xyz)$. In this scenario, $p(abc|xyz)$ is said to be bilocal if it can be written as
$p(abc|xyz)=\int\int d\lambda d\mu p_{1}(\lambda)p_{2}(\mu)p(a|x,\lambda)p(b|y,\lambda,\mu)p(c|z,\mu),$
where $\lambda$ and $\mu$ characterize the hidden variables of the systems produced by the sources $\rho_{AB}$ and $\rho_{BC}$, respectively [65, 84]. Otherwise, it is called non-bilocal.
In order to detect non-bilocal correlations generated by the network, it is often necessary to find suitable measurements that violate the following bilocal inequality
$\mathcal{S}\equiv\sqrt{|I|}+\sqrt{|J|}\leq 2,$ (2.1)
whose maximum quantum violation is $2\sqrt{2}$ and is attainable. Here
$I\equiv\sum_{x,z}\langle A_{x}B_{0}C_{z}\rangle=\langle(A_{0}+A_{1})B_{0}(C_{0}+C_{1})\rangle,$
$J\equiv\sum_{x,z}(-1)^{x+z}\langle A_{x}B_{1}C_{z}\rangle=\langle(A_{0}-A_{1})B_{1}(C_{0}-C_{1})\rangle$
as introduced in Ref. [63]:
$\langle A_{x}B_{y}C_{z}\rangle$ $=$ $\sum_{a,b,c=0}^{1}(-1)^{a+b+c}{\rm tr}((A_{a|x}B_{b|y}C_{x|z})\rho_{AB}\otimes\rho_{BC})$
$=$ $\sum_{a,b,c=0}^{1}(-1)^{a+b+c}p(abc|xyz).$
Here $A_{x}=\sum_{a}(-1)^{a}A_{a|x}$, $B_{y}=\sum_{b}(-1)^{b}B_{b|y}$ and $C_{z}=\sum_{z}(-1)^{z}C_{c|z}$, where $A_{a|x},\ B_{b|y}$, and $C_{c|z}$ are the positive operator-valued measurements (POVMs) performed by Alice, Bob and Charles, respectively.
Mutually-commuting von Neumann algebra models. In QFT, the observables for Alice, Bob, and Charles are associated with three mutually-commuting von Neumann algebras $\mathcal{M}_{\mathcal{A}}$, $\mathcal{M}_{\mathcal{B}}$, $\mathcal{M}_{\mathcal{C}}$. Therefore, our model encompasses both the non-relativistic quantum mechanics scenario and the quantum field theory scenario. The idea of this model is similar to that in Refs. [81, 83].
Definition 2.1.
(Ternary Mutually-commuting von Neumann Algebra Models of Tripartite Quantum Systems) Let $\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}}$ be von Neumann subalgebras of $\mathcal{B}(\mathcal{H})$ over some Hilbert space $\mathcal{H}$, which are mutually commuting, i.e., $\mathcal{M}_{i}\subset\mathcal{M}_{j}^{\prime}$ with $i\neq j\in\{\mathcal{A},\mathcal{B},\mathcal{C}\}$, where $\mathcal{M}_{j}^{\prime}$ is the communtant of $\mathcal{M}_{j}$. The generated von Neumann algebra
$\mathcal{M}_{\mathcal{ABC}}=(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}}\vee\mathcal{M}_{\mathcal{C}})^{\prime\prime}.$
We refer to the above model as the TMCvNA model. When $\mathcal{M}_{\mathcal{ABC}}\simeq\mathcal{M}_{\mathcal{A}}\otimes\mathcal{M}_{\mathcal{B}}\otimes\mathcal{M}_{\mathcal{C}}$, it is called the tensor product algebra model. In this paper, for any $A\in\mathcal{M}_{\mathcal{A}}$, $B\in\mathcal{M}_{\mathcal{B}}$, $C\in\mathcal{M}_{\mathcal{C}}$, we always assume that they are Hermitian.
We intend to use the above model of ternary mutually commuting von Neumann algebras to describe the entanglement swapping network in Fig. 1. We note that there is no correlation between the parties Alice and Charles in the network. Mathematically, this independence can be described by the following formula: a network state $\tau$ of the entanglement swapping network should be a state in $\mathcal{M}_{\mathcal{ABC}}^{*}$, the dual space of $\mathcal{M}_{\mathcal{ABC}}$, satisfying
$\tau(AC)=\tau(A)\tau(C)$ ()
for any $A\in\mathcal{M}_{\mathcal{A}},C\in\mathcal{M}_{\mathcal{C}}$. We call (*) the independent condition. This assumption will be used throughout this paper.
Definition 2.2.
The ternary mutually-commuting von Neumann algebra model of entanglement swapping networks is the ternary mutually-commuting von Neumann algebra model of tripartite quantum systems with all states satisfying the independence condition (*).
3 Bilocal inequalities and their bounds
In this section, we further analyze the conditions under which the bilocal inequality holds or is violated in the TMCvNA model of an entanglement swapping network. Specifically, in this model, we can construct the bilocal inequality analogous to that in the non-relativistic setting. Let
$I_{\tau}=\tau\bigl((A_{0}+A_{1})B_{0}(C_{0}+C_{1})\bigr),$
$J_{\tau}=\tau\bigl((A_{0}-A_{1})B_{1}(C_{0}-C_{1})\bigr),$
where $\tau$ is a state on $\mathcal{M}_{\mathcal{ABC}}$ satisfying the independent condition (*). Here, $\tau\bigl(A_{x}B_{y}C_{z}\bigr)=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}\tau({A}_{a|x}{B}_{b|y}{C}_{c|z})$ and $A_{x}=\sum_{a}(-1)^{a}A_{a|x}$, $B_{y}=\sum_{b}(-1)^{b}B_{b|y}$, $C_{z}=\sum_{z}(-1)^{z}C_{c|z}$, where $A_{a|x}$, $B_{b|y}$, and $C_{c|z}$ are the POVMs performed by Alice, Bob, and Charles, respectively. Moreover, the network correlation $\hat{p}=p(\alpha\beta\gamma|xyz)$ in the TMCvNA model is defined as
$p(\alpha\beta\gamma|xyz)=\tau({A}_{\alpha|x}{B}_{\beta|y}{C}_{\gamma|z}),$
In the TMCvNA model, analogous to Ineq. (2.1), we set
$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}.$ (3.1)
We say that the state $\tau$ together with the observables $A_{x}$, $B_{y}$, $C_{z}$ satisfies the bilocal inequality if $\mathcal{S}_{\tau}\leq 2$, and violates it if $\mathcal{S}_{\tau}>2$.
The following conclusion indicates that in the TMCvNA of entanglement swapping networks, the supremum of $\mathcal{S}_{\tau}$ defined in Eq. (3.1) is $2\sqrt{2}$. This coincides with the case in non-relativistic quantum mechanics, where the bilocal quantity $\mathcal{S}$ in Ineq. (2.1) attains a maximal violation of $2\sqrt{2}$ allowed by quantum resources.
Theorem 3.1.
For the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, we always have $\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2\sqrt{2}.$
Proof According to the Gelfand-Namark-Segal (GNS) construction, there is a -representation $\pi_{\tau}:\mathcal{M}_{\mathcal{ABC}}\rightarrow\mathcal{B}(\mathcal{H}_{\tau})$ and a cyclic vector $\Omega\in\mathcal{H}_{\tau}$ such that the set $\{\pi_{\tau}(O)\Omega:\ O\in\mathcal{M}_{\mathcal{ABC}}\}$ is dense in $\mathcal{H}_{\tau}$. It follows by applying the Cauchy-Schwarz inequality that
$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}$
$=\sqrt{|\tau((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\tau((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}$
$\leq\sqrt{2}\sqrt{|\tau(B_{0}(A_{0}+A_{1})(C_{0}+C_{1}))|+|\tau(B_{1}(A_{0}-A_{1})(C_{0}-C_{1}))|}$
$=\sqrt{2}\sqrt{|\langle\pi_{\tau}(B_{0})\Omega,\pi_{\tau}((A_{0}+A_{1})(C_{0}+C_{1}))\Omega\rangle|+|\langle\pi_{\tau}(B_{1})\Omega,\pi_{\tau}((A_{0}-A_{1})(C_{0}-C_{1}))\Omega\rangle|}$
$\leq\sqrt{2}\sqrt{\|\pi_{\tau}(B_{0})\Omega\|\|\pi_{\tau}((A_{0}+A_{1})(C_{0}+C_{1}))\Omega\|+\|\pi_{\tau}(B_{1})\Omega\|\|\pi_{\tau}((A_{0}-A_{1})(C_{0}-C_{1}))\Omega\|}$
$\leq\sqrt{2}\sqrt{\sqrt{\tau((A_{0}+A_{1})^{2}(C_{0}+C_{1})^{2})}+\sqrt{\tau((A_{0}-A_{1})^{2}(C_{0}-C_{1})^{2})}}$
$\leq\sqrt{2}\sqrt{\sqrt{\tau(A_{0}+A_{1})^{2}+\tau(A_{0}-A_{1})^{2}}\sqrt{\tau(C_{0}+C_{1})^{2}+\tau(C_{0}-C_{1})^{2}}}$
$=\sqrt{2}\sqrt{\sqrt{2\tau(A_{0}^{2}+A_{1}^{2})}\sqrt{2\tau(C_{0}^{2}+C_{1}^{2})}}$
$\leq\sqrt{2}\sqrt{2\sqrt{2}\sqrt{2}}$
$=2\sqrt{2}.$
The final inequality invokes the condition that $I\leq A_{i}\leq I,\ -I\leq C_{i}\leq I$ and the positivity property of $\tau$. $\Box$
Building on the results above, we now investigate how the quantity $\mathcal{S}_{\tau}$ in Eq. (3.1) depends on the abelianness of the algebras $\mathcal{M}_{\mathcal{A}}$, $\mathcal{M}_{\mathcal{B}}$, and $\mathcal{M}_{\mathcal{C}}$. Specifically, the results indicate that in entanglement swapping networks, the abelianness of the three algebras plays distinct roles in reducing the upper bound of the inequality to 2, i.e., determining the conditions under which no violation of the bilocal inequality can occur. This is not a simple generalization of the bipartite Bell scenario [47], where, with only two systems, Summers et al. showed that if one of these two algebras is abelian, the upper bound of the Bell inequality is 2.
Theorem 3.2.
For the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, then
$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2.$
Proof Since $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, respectively, the eight elements
$A_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{0}A_{0})(1+\epsilon_{1}A_{1}),\ \ \ C_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{1}C_{0})(1+\epsilon_{1}C_{1})$
with $\epsilon_{0},\epsilon_{1}\in\{+,-\}$ are positive. By direct computation, one obtains that
$A_{0}+A_{1}=2(A_{++}-A_{--}),\ C_{0}+C_{1}=2(C_{++}-C_{--}),$
$A_{0}-A_{1}=2(A_{+-}-A_{-+}),\ C_{0}-C_{1}=2(C_{+-}-C_{-+}).$
So one obtains that
$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\tau((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}$
$=2\left(\sqrt{|\tau((A_{++}-A_{--})B_{0}(C_{++}-C_{--}))|}+\sqrt{|\tau((A_{+-}-A_{-+})B_{1}(C_{+-}-C_{-+}))|}\right)$
$=2\left(\sqrt{|\tau(A_{++}B_{0}C_{++})-\tau(A_{++}B_{0}C_{--})-\tau(A_{--}B_{0}C_{++})+\tau(A_{--}B_{0}C_{--})|}\right$.
$\qquad\left.+\sqrt{|\tau(A_{+-}B_{1}C_{+-})-\tau(A_{+-}B_{1}C_{-+})-\tau(A_{-+}B_{1}C_{+-})+\tau(A_{-+}B_{1}C_{-+})|}\right)$
$\leq 2\left(\sqrt{|\tau(A_{++}C_{++})+\tau(A_{++}C_{--})+\tau(A_{--}C_{++})+\tau(A_{--}C_{--})|}\right$.
$\qquad\left.+\sqrt{|\tau(A_{+-}C_{+-})+\tau(A_{+-}C_{-+})+\tau(A_{-+}C_{+-})+\tau(A_{-+}C_{-+})|}\right)$
$=2\left(\sqrt{|\tau(A_{++}+A_{--})||\tau(C_{++}+C_{--})|}+\sqrt{|\tau(A_{+-}+A_{-+})||\tau(C_{+-}+C_{-+})|}\right)$
$\leq 2\sqrt{\tau(A_{++}+A_{--})+\tau(A_{+-}+A_{-+})}\sqrt{\tau(C_{++}+C_{--})+\tau(C_{+-}C_{-+})}$
$=2\sqrt{1}\sqrt{1}=2,$
where the first inequality follows the fact that $I\leq B_{i}\leq I\ (i=0,1)$ and the order-preserving of state $\tau$, and the second inequality holds because of the Cauchy-Schwarz inequality and the non-negativeness of $A_{\epsilon_{0}\epsilon_{1}}\ \text{and}\ \ C_{\epsilon_{0}\epsilon_{1}}$. $\Box$
The above theorem illustrates a phenomenon: the violation of the Bell-type inequality, i.e., $2<\mathcal{S}_{\tau}\leq 2\sqrt{2}$ can serve as an indicator of the non-abelianness of the underlying algebras. Applying the theorem, we can infer from a violation of the inequality in Theorem 3.2 that at least one of the algebras $\mathcal{M_{A}}$ and $\mathcal{M_{C}}$ is non-abelian.
In the following, we define a quantity
$\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}).$
Combining Theorems 3.1 and 3.2, one naturally obtains the following corollary.
Corollary 3.3.
In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks,
(1) for any state $\tau$ and any choice of observables $A_{x}\in\mathcal{M}_{\mathcal{A}},\ B_{y}\in\mathcal{M}_{\mathcal{B}},\ C_{z}\in\mathcal{M}_{\mathcal{C}}$ in a scheme with two inputs and two outputs, the quantity $\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})$ satisfies
$2\leq\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})\leq 2\sqrt{2}.$
(2) if $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, then $\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=2$.
(3) for any states $\phi,\psi\in[(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}}\vee\mathcal{M}_{\mathcal{C}})^{\prime\prime}]^{*}$, the following inequality holds:
$|\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})-\mathcal{S}(\psi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})|\leq k\sqrt{\|\phi-\psi\|},$
where $k$ is a positive constant. Consequently, the functional $\phi\mapsto\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|})$ is norm continuous.
Proof To show (1), note that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and combining this with the proof of Theorem 3.1, we obtain (1).
(2) holds by the fact that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and by Theorem 3.2.
To prove (3), note that by the representations of $I_{\tau}$ and $J_{\tau}$, together with the facts that $|\sup x-\sup y|\leq\sup|x-y|$ and the triangle inequality, one can obtain
$|\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})-\mathcal{S}(\psi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})|$
$=\left|\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|}\right)-\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{|I_{\psi}|}+\sqrt{|J_{\psi}|}\right)\right|$
$=\left|\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{|\phi((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\phi((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}\right)\right$.
$\qquad\left.-\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{|\psi((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\psi((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}\right)\right|$
$\leq\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left|\sqrt{|\phi((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\phi((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}\right$.
$\qquad\left.-\sqrt{|\psi((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}-\sqrt{|\psi((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}\right|$
$\leq\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\left|\sqrt{|\phi((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}-\sqrt{|\psi((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}\right|\right$.
$\qquad\left.+\left|\sqrt{|\phi((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}-\sqrt{|\psi((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}\right|\right)$
$\leq\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{|(\phi-\psi)((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|(\phi-\psi)((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}\right)$
$\leq\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{\|\phi-\psi\|\,\|(A_{0}+A_{1})B_{0}(C_{0}+C_{1})\|}+\sqrt{\|\phi-\psi\|\,\|(A_{0}-A_{1})B_{1}(C_{0}-C_{1})\|}\right)$
$\leq k\sqrt{\|\phi-\psi\|},$
where the fourth one follows the Cauchy Schwarz inequality, and the last obeys the norm for elements of $\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}}$ are bounded. So $\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|})$ is norm continuous in the state $\phi$. $\Box$
4 Maximal violation of bilocal inequalities and algebraic structures
In this section, we aim to show that the violation of bilocal inequalities, in particular the maximal violation, can reflect the structural properties of the algebra. Here, violation refers to exceeding the bound for $\mathcal{S}_{\tau}$ in Eq. (3.1), while maximal violation means attaining the value $2\sqrt{2}$ for the same quantity.
The following theorem analyzes the conditions for maximal violation.
Theorem 4.1.
For the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if the network state $\tau\in\mathcal{M}_{\mathcal{ABC}}^{*}$ is faithful, then
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$
if and only if $\tau(A_{i}^{2}A)=\tau(A),\ \tau(B_{i}^{2}B)=\tau(B),\ \tau(C_{i}^{2}C)=\tau(C)$, and $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=0,\ \tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$ for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$ with $i\in\{0,1\}$.
Proof If $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$, it follows from the proof of Theorem 3.1 that for any $t\in[0,1]$, these equalities hold:
$|\tau(B_{0}(A_{0}+A_{1})(C_{0}+C_{1}))|=|\tau(B_{1}(A_{0}-A_{1})(C_{0}-C_{1}))|,$ (4.1)
$\pi_{\tau}(B_{0})\Omega=k_{0}\pi_{\tau}[(A_{0}+A_{1})(C_{0}+C_{1})]\Omega,$ (4.2)
$\pi_{\tau}(B_{1})\Omega=k_{1}\pi_{\tau}[(A_{0}-A_{1})(C_{0}-C_{1})]\Omega,$ (4.3)
$\|\pi_{\tau}(B_{0})\Omega\|=\|\pi_{\tau}(B_{1})\Omega\|=1,$ (4.4)
$\tau[(A_{0}+A_{1})^{2}]=t\tau[(C_{0}+C_{1})^{2}],$ (4.5)
$\tau[(A_{0}-A_{1})^{2}]=t\tau[(C_{0}-C_{1})^{2}],$ (4.6)
$A_{0}^{2}+A_{1}^{2}=2I,C_{0}^{2}+C_{1}^{2}=2I.$ (4.7)
From (4.7), one gets $A_{i}^{2}=I,\ C_{i}^{2}=I\ (i=0,1)$ because $I\leq A_{i},C_{i}\leq I$, therefore $\tau(A_{i}^{2}A)=\tau(A),\ \tau(C_{i}^{2}C)=\tau(C)$ for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$.
Now, let us show the proof for $\tau(B_{i}^{2}B)=\tau(B),\ \tau[(A_{0}A_{1}+A_{1}A_{0})A]=0$, and $\tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$ for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$, which implies $B_{i}^{2}=I,$ $A_{0}A_{1}+A_{1}A_{0}=0,$ and $C_{0}C_{1}+C_{1}C_{0}=0$. According to Eqs. (4.5)-(4.7), i.e.,
$\left\{\begin{array}[]{c}\tau(A_{0}^{2}+A_{1}^{2}+A_{0}A_{1}+A_{1}A_{0})=t\tau(C_{0}^{2}+C_{1}^{2}+C_{0}C_{1}+C_{1}C_{0})\\ \tau(A_{0}^{2}+A_{1}^{2}-A_{0}A_{1}-A_{1}A_{0})=t\tau(C_{0}^{2}+C_{1}^{2}-C_{0}C_{1}-C_{1}C_{0})\\ A_{0}^{2}+A_{1}^{2}=2I,\ C_{0}^{2}+C_{1}^{2}=2I,\end{array}\right$.
we can get
$t=1,\ \tau(A_{0}A_{1}+A_{1}A_{0})=\tau(C_{0}C_{1}+C_{1}C_{0}).$
Then combing conditions (4.2), (4.3) and condition (4.4), one can get
$k_{0}^{2}\tau[(A_{0}+A_{1})^{2}(C_{0}+C_{1})^{2}]=1,\ k_{1}^{2}\tau[(A_{0}-A_{1})^{2}(C_{0}-C_{1})^{2}]=1,$
i.e.,
$|k_{0}|=\frac{1}{2+\tau(X)},\ \ \ |k_{1}|=\frac{1}{2-\tau(X)}$ (4.8)
because of $\tau(X)=\tau(Y)$, where $X=A_{0}A_{1}+A_{1}A_{0},\ Y=C_{0}C_{1}+C_{1}C_{0}$. Then according to Eqs. (4.2), (4.3), and (4.1),
$|\tau[B_{0}(A_{0}+A_{1})(C_{0}+C_{1})]|=|k_{0}||\tau[(A_{0}+A_{1})^{2}(C_{0}+C_{1})^{2}]|,$
$|\tau[B_{1}(A_{0}-A_{1})(C_{0}-C_{1})]|=|k_{1}||\tau[(A_{0}-A_{1})^{2}(C_{0}-C_{1})^{2}]|.$
Substituting (4.8) and (4.1) to the above equations, we have
$\frac{1}{2+\tau(X)}(2+\tau(X))^{2}=\frac{1}{2-\tau(X)}(2-\tau(X))^{2}$
deriving $\tau(X)=\tau(Y)=0$. It follows from $\tau(X)=0=\tau(Y)$ and $\|\pi_{\tau}(B_{i})\Omega\|=1$ that
$k_{0}^{2}=k_{1}^{2}=\frac{1}{4}$
from Eq. (4.8). Since
$|\langle\Omega,\pi_{\tau}(B_{0}^{2})\Omega\rangle|$ $=$ $|\frac{1}{4}\langle\Omega,\pi_{\tau}[(A_{0}+A_{1})^{2}(C_{0}+C_{1})^{2}]\Omega\rangle|$
$\leq$ $\frac{1}{4}\sqrt{\|\pi_{\tau}[(A_{0}+A_{1})(C_{0}+C_{1})]\Omega\|}\sqrt{\|\pi_{\tau}[(A_{0}+A_{1})(C_{0}+C_{1})]\Omega\|}$
$=$ $\frac{1}{4}\tau[(A_{0}+A_{1})^{2}]\tau[(C_{0}+C_{1})^{2}]=1,$
and $\tau(B_{0}^{2})=|\langle\Omega,\pi_{\tau}(B_{0}^{2})\Omega\rangle|=1$, this implies that $\pi_{\tau}[(B_{0}^{2})]\Omega=\Omega$, so
$\tau(B_{0}^{2}B)=\langle\Omega,\pi_{\tau}(B_{0}^{2})\pi_{\tau}(B)\Omega\rangle=\langle\pi_{\tau}(B_{0}^{2})\Omega,\pi_{\tau}(B)\Omega\rangle=\langle\Omega,\pi_{\tau}(B)\Omega\rangle=\tau(B).$
Similarly $\tau(B_{1}^{2}B)=\tau(B)$ for any $B\in\mathcal{M}_{\mathcal{B}}.$ Furthermore, it follows from Eq. (4.2) and $\pi_{\tau}[(B_{0}^{2})]\Omega=\Omega$ that
$\pi_{\tau}(2X+2Y+XY)\Omega=0.$
So $\tau[(2X+2Y+XY)^{2}]=0,$ implying that
$\tau(X^{2})=\tau(Y^{2})=0.$
Combining the faithfulness, non-negativity of the state $\tau$ and the self-adjointness of X,YX,\ Y. Then $X=Y=0$, i.e.,
$A_{0}A_{1}+A_{1}A_{0}=C_{0}C_{1}+C_{1}C_{0}=0,$
and implies that $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=\tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$ for any $A\in\mathcal{M}_{\mathcal{A}},\ C\in\mathcal{M}_{\mathcal{C}}$.
It is straightforward to prove the converse process, as we check that Eqs. (4.1)-(4.7) hold if $\tau(A_{i}^{2}A)=\tau(A),\ \tau(B_{i}^{2}B)=\tau(B),\ \tau(C_{i}^{2}C)=\tau(C)$, and $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=0,\ \tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$. We complete the proof. $\Box$
To further elucidate the algebraic relations presented in Theorem 4.1, we provide the following corollary.
Corollary 4.2.
For the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, the bilocal inequality can be maximally violated if and only if $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ contain subalgebras isomorphic to $M_{2}(\mathbb{C})$ and there exists a faithful state $\tau\in\mathcal{M}_{\mathcal{ABC}}^{*}$ that satisfies the independent condition (*): $\tau(AC)=\tau(A)\tau(C)$ for all $A\in\mathcal{M}_{\mathcal{A}},\ C\in\mathcal{M}_{\mathcal{C}}$.
Proof Note that for any von Neumann algebra, there always exists a faithful state $\tau\in\mathcal{M}_{\mathcal{ABC}}^{*}$.
$(\Leftarrow)$ Suppose $\mathcal{M}_{\mathcal{A}}$ (resp. $\mathcal{M}_{\mathcal{C}}$) contains a subalgebra $\mathcal{M}_{\mathcal{A}}^{sub}\simeq M_{2}(\mathbb{C})$ (resp. $\mathcal{M}_{\mathcal{C}}^{sub}\simeq M_{2}(\mathbb{C})$) and $\tau$ satisfies ()(*).
Then there exist operators $A_{0},A_{1},$ and $A_{2}:=-\frac{{\rm{i}}}{2}[A_{0},A_{1}]$ in $\mathcal{M}_{\mathcal{A}}^{sub}$ such that they anticommute and $A_{i}^{2}=I$ for $i\in\{0,1,2\}$, where ${\rm{i}}^{2}=-1$. Consequently, we obtain
$\tau(A_{i}^{2}A)=\tau(A),\ \tau[(A_{0}A_{1}+A_{1}A_{0})A]=0$
for any $A\in\mathcal{M}_{\mathcal{A}}$. Similarly for $C\in\mathcal{M}_{\mathcal{C}}$ with $C_{0},C_{1},C_{2}:=-\frac{{\rm{i}}}{2}[C_{0},C_{1}]$.
Setting $B_{0}=B_{1}=I$ gives $\tau(B_{i}^{2}B)=\tau(B)$ for all $B\in\mathcal{M}_{\mathcal{B}}$. By Theorem 4.1, these operators yield the maximal violation $2\sqrt{2}$ of the quantity $\mathcal{S}_{\tau}$ in Eq. (3.1).
$(\Rightarrow)$ Conversely, assume $\mathcal{S}_{\tau}$ in Eq. (3.1) attains the maximal violation $2\sqrt{2}$.
Then for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$ and $i\in\{0,1\}$, we have $\tau(A_{i}^{2}A)=\tau(A),\tau(B_{i}^{2}B)=\tau(B),\tau(C_{i}^{2})=\tau(C)$, and $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=0,\ \tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$.
Taking $A=A_{0}A_{1}+A_{1}A_{0}$ and using the faithfulness of state $\tau$ together with $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=0$, one gets
$A_{0}A_{1}+A_{1}A_{0}=0,$
i.e., $A_{0}A_{1}=-A_{1}A_{0}$. The algebra generated by $A_{0},A_{1}$ is
$\mathfrak{A}(A_{0},A_{1}):=\{\sum_{k}\alpha_{k}A_{0}^{m}A_{1}^{n}|\alpha_{k}\in\mathbb{C},m,n\in\mathbb{N}\}.$
From $\tau(A_{i}^{2}A)=\tau(A)$ and $I\leq A_{i}\leq I\ (i\in\{0,1\})$, setting $A=I$ gives $\tau(I-A_{i}^{2})=0$. Faithfulness of state $\tau$ then implies $A_{i}^{2}=I$. The same reasoning yields $B_{i}^{2}=I\ \textmd{and}\ C_{i}^{2}=I$.
Because $A_{i}^{2}=I$, one gets $\sum_{k}\alpha_{k}A_{0}^{m}A_{1}^{n}=\alpha_{0}I+\alpha_{1}A_{0}+\alpha_{2}A_{1}+\alpha_{3}A_{0}A_{1}$. Hence,
$\mathfrak{A}(A_{0},A_{1})={\rm span}\{I,A_{0},A_{1},-\frac{{\rm{i}}}{2}[A_{0},A_{1}]\}\simeq M_{2}(\mathbb{C}).$
Analogously, $\mathfrak{A}(C_{0},C_{1})\simeq M_{2}(\mathbb{C})$. So this proof is completed. $\Box$
The above results show a close connection between the maximal violation of Bell-type inequalities and algebraic structures. This connection can be more clearly demonstrated when all algebras are factors.
Corollary 4.3.
For the ternary tensor product algebra models of entanglement swapping networks, i.e., $\mathcal{M}_{\mathcal{ABC}}=\mathcal{M}_{\mathcal{A}}\otimes\mathcal{M}_{\mathcal{B}}\otimes\mathcal{M}_{\mathcal{C}}$, assume that $\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}}$ are factors. Then the bilocal inequality cannot be maximally violated if and only if either $\mathcal{M}_{\mathcal{A}}$ or $\mathcal{M}_{\mathcal{C}}$ is type I2k+1 for some finite positive integer $k$.
5 Discussion and Conclusions
We investigate Bell type inequality for entanglement swapping network in von Neumann algebraic framework, extending the paradigm that Bell violation from observable algebra structure (notably type III factors). We identify algebraic constraints governing inequality violation, linking network nonlocality to the noncommutative structure of the underlying algebras, and further show that maximal violation conditions can reverse-engineer the type classication of von Neumann algebras.
This work represents merely the beginning of a much broader inquiry. Our current model focuses primarily on the simplest nontrivial network: the entanglement swapping scenario with two independent sources. The generalization of these results to arbitrary multipartite quantum networks remains an important open problem. In more complex architectures, the interplay between multiple independent sources and the commutation relations of their associated algebras is expected to reveal even richer structures of nonlocality in networks represented by mutually-commuting von Neumann algebras.
Conflict of Interest  The authors declare no conflict of interest.
Acknowledgements.
We thank the referees for their time and comments.